### v2

# On computability of equilibrium states

Abstract.
Equilibrium states are natural dynamical analogs of Gibbs states in thermodynamic formalism. This paper is devoted to the study of their computability in the sense of Computable Analysis. We show that the unique equilibrium state associated to a pair of a computable, topologically exact, distance-expanding, open map $T\colon X\rightarrow X$ and a computable Hölder continuous potential $\varphi\colon X\rightarrow\mathbb{R}$ is always computable. Furthermore, the Hausdorff dimension of the Julia set and the equilibrium state for the geometric potential of a computable hyperbolic rational map are computable. On the other hand, we introduce a mechanism to establish non-uniqueness of equilibrium states. We also present some computable dynamical systems whose equilibrium states are all non-computable.
Key words and phrases:
computability, computable analysis, equilibrium state, thermodynamic formalism, uniformly expanding, hyperbolic rational map.
2020 Mathematics Subject Classification:
Primary: 03D78; Secondary: 37D35, 37D20, 37F15
1. Introduction
The decidability problem was the third question of Hilbert’s program proposed by Hilbert in Paris in 1900. In the 1930s, the Halting Problem given by Turing provided a counterexample to this question. Another significant contribution of Turing was the construction of a definite method of computations called Turing machine. Turing machines give a standard model of computation in discrete settings and have become the foundation of modern digital computers. Nowadays, however, most digital computers are used for computations in continuous settings. The research on computations with real numbers began with Turing’s work on the original definition of computable real numbers in 1937 ([Tu37]). The work of Banach and Mazur in 1937 ([BM37]) provided definitions of computability for real objects (such as subsets of $\mathbb{R}^{n}$ and functions $f\colon\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$). The “bit model” has been developed, serving as a foundation of the tradition of Computable Analysis ([Gr55, La55, Ko91, Weih00]). Computable Analysis on dynamical systems. In recent decades, there has been dramatic growth in research on the computability and computational complexity of many objects generated by dynamical systems (see for example, [Ya21] and references therein). Initial value sensitivity and typical instability of many interesting systems imply that what is being observed on the computer screen could be completely unrelated to what was meant to be simulated. This paper is devoted to the study of computability questions for a family of invariant measures called equilibrium states (see Definition 2.4). It expands and continues the line of inquiry initiated by [BBRY11] of the first-named author and Braverman, Rojas, and Yampolsky. Recall that for a rational map $f(z)=P(z)/{Q(z)}$ on the Riemann sphere, where $P(z)$ and $Q(z)$ are mutually prime polynomials, with degree $\deg(f)=\max\{\deg(P),\,\deg(Q)\}\geqslant 2$, the Julia set $\mathcal{J}_{f}$ is defined as the locus of the chaotic dynamics of $f$, namely, the complement of the set where the dynamics of $f$ is stable. Moreover, the $*{}^{*}$-limit of the average of Dirac measures supported on the preimage points of a repelling periodic point $w$ is an $f$-invariant measure $\lambda$ (independent of the choice of $w$) called the Brolin–Lyubich measure. In [BBRY11], a uniform machine was designed to compute the Brolin–Lyubich measure for any given rational map. Together with the existence of polynomials whose coefficients are all computable, but whose Julia sets are non-computable from the groundbreaking works of Braverman and Yampolsky ([BY06, BY09]) dating back to a question of Milnor [BY06, Section 1], the computability of Brolin–Lyubich measures leads to a conflict: heuristically speaking, a measure contains more information than its support, but in Computable Analysis, there exists a computable invariant probability measure whose support is, however, non-computable. Such a conflict can be reconciled by considering these two results as the computability properties of the same physical objects from the geometric perspective and the statistical perspective, respectively. In this era of artificial intelligence and data science, the latter undoubtedly deserves closer investigations. Thermodynamic formalism is a powerful method for creating invariant measures with prescribed local behavior under iterations of the dynamical system. This theory, inspired by statistical mechanics, was created by Ruelle, Sinai, and others in the seventies ([Dob68, Si72, Bo75, Wa82]). Since then, thermodynamic formalism has been applied in many classical contexts (see for example, [Bo75, Ru89, Pr90, KH95, Zi96, MauU03, BS03, Ol03, Yu03, PU10, MayU10]) and has remained at the frontier of researches in dynamical systems. Among many other applications, thermodynamic formalism has, since its very early days, played a central role in the study of statistical properties of dynamical systems. The key objects of investigation in thermodynamic formalism are invariant measures called equilibrium states. More precisely, for a continuous map $T\colon X\rightarrow X$ of a compact metric space $(X,\rho)$, a $T$-invariant Borel probability measure $\mu$ on $X$, and a real-valued continuous function $\varphi\colon X\rightarrow\mathbb{R}$, called potential, one can define the measure-theoretic pressure and the topological pressure to describe the degree of complexity of the given dynamical system from different perspectives. The Variational Principle implies that the supremum over all $T$-invariant Borel probability measures on $X$ of the measure-theoretic pressure is equal to the topological pressure. We say that a Borel probability measure $\mu$ is an equilibrium state of $T$ for $\varphi$ if $\mu$ maximizes the measure-theoretic pressure. In particular, if $\varphi$ is a constant function, then an equilibrium state reduces to a measure of maximal entropy (see Subsection 2.2). In particular, the Brolin–Lyubich measure of a rational map $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is the measure of maximal entropy of $f$. Statement of main results. The main theorems of this paper include two positive results on the computability of equilibrium states in uniformly expanding systems, an application to complex dynamics, an application to the classical theory of thermodynamic formalism (independent of the interest from Computable Analysis) on non-uniqueness of equilibrium states, and some counterexamples. Computability of equilibrium states. Our first result concerns the computability of equilibrium states. Here is a list of assumptions used in this context.
Assumption A.
  1. (i)
    $(X,\,\rho,\,\mathcal{S})$ is a computable metric space and $X$ is a recursively compact set (in the sense of Definition 3.14).
  2. (ii)
    $\varphi\colon X\rightarrow\mathbb{R}$ is a real-valued Hölder continuous potential with some exponent $v_{0}$ and constant $a_{0}$ (in the sense of Definition 2.6).
  3. (iii)
    $T\colon X\rightarrow X$ is an open, topologically exact, continuous map that is distance-expanding (in the sense of Definition 2.1) with respect to the metric $\rho$ with some constants $\eta$, $\lambda$, and $\xi$.
  4. (iv)
    $\varphi\colon X\rightarrow\mathbb{R}$ and $T\colon X\rightarrow X$ are both computable (in the sense of Definition 3.9).
By classical results (see for example, Proposition 2.5 and [PU10, Corollary 5.2.14]), the Assumption A (i) through (iii) imply that there exists a unique equilibrium state of $T$ for $\varphi$. The Assumption A (iv) allows us to input the information of the system in the computer. For systems satisfying the above assumptions, we prove the following theorem.
Theorem A.
Let the quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfy the Assumption A in Section 1, and $\mu$ be the unique equilibrium state of the map $T$ for the potential $\varphi$. Then $\mu$ is computable.
For a more precise version of Theorem A, see Theorem 1.1 below. The Assumption A requires that the behavior of the dynamical system has nice properties in the whole space. However, generally speaking, the dynamical behavior may be merely good enough in an invariant subspace rather than in the whole space. Hence, we prove Theorem B, a “subspace” version of Theorem A. We give a list of assumptions that describes the systems considered in Theorem B as follows.
Assumption B.
  1. (i)
    $(X,\,\rho,\,\mathcal{S})$ is a computable metric space.
  2. (ii)
    $(D,\,\rho,\,\mathcal{S}^{\prime})$ is a computable metric space, where $\mathcal{S}^{\prime}$ is a uniformly computable sequence of points in the computable metric space $(X,\,\rho,\,\mathcal{S})$. $D$ is a recursively compact set in the computable metric space $(D,\,\rho,\,\mathcal{S}^{\prime})$.
  3. (iii)
    $\varphi\colon D\rightarrow\mathbb{R}$ is a real-valued Hölder continuous potential with some exponent $v_{0}$ and constant $a_{0}$ (in the sense of Definition 2.6).
  4. (iv)
    $T\colon X\rightarrow X$ is a map such that $D$ is completely $T$-invariant. $T|_{D}\colon D\rightarrow D$ is an open, topologically exact, continuous map that is distance-expanding with respect to the metric $\rho$ with some constants $\eta$, $\lambda$, and $\xi$.
  5. (v)
    $\varphi\colon D\rightarrow\mathbb{R}$ and $T|_{D}\colon D\rightarrow\mathbb{R}$ are both computable in the computable metric space $(D,\,\rho,\,\mathcal{S}^{\prime})$.
The Assumption B (ii) through (iv) give the uniqueness of the equilibrium state of $T|_{D}$ for the potential $\varphi$. Together with Theorem 1.1, the Assumption B (ii) through (v) imply that the equilibrium state is computable in the computable metric space $(D,\,\rho,\,\mathcal{S}^{\prime})$. Then the Assumption B (ii) and (iii) allow us to deduce the computability in the computable metric space $(X,\,\rho,\,\mathcal{S})$ from the computability in the computable metric space $(D,\,\rho,\,\mathcal{S}^{\prime})$. Therefore, for the systems satisfying the above assumptions, we prove the following theorem.
Theorem B.
Let the septet $(X,\,D,\,\rho,\,\mathcal{S},\,\mathcal{S}^{\prime},\,\varphi,\,T)$ satisfy the Assumption B in Section 1, and $\mu$ be the unique equilibrium state of $T|_{D}$ for the potential $\varphi$. Then $\mu$ is computable.
For a more precise version of Theorem B, see Theorem 1.2 below. As an application, we apply Theorem 1.2 to study the equilibrium states of hyperbolic rational maps on the Riemann sphere associated to the geometric potentials. We say that a rational map is hyperbolic if it is distance-expanding after iterations on the Julia set with respect to the spherical metric. Hyperbolic rational maps are abundant and are expected to be dense among all the rational maps. In the case of quadratic rational maps, see the renowned MLC Conjecture (refer to [Mil06, Appendix G] for reference). The studies on equilibrium states as well as the theory of thermodynamic formalism for hyperbolic rational maps are by now well-developed and established. Given $t\in\mathbb{R}$, it is well-known that a hyperbolic rational map $f$ restricted to the Julia set $\mathcal{J}_{f}$ admits a unique non-atomic equilibrium state $\mu_{t,f}$ which is supported on $\mathcal{J}_{f}$ for the geometric potential $t\varphi_{f}(z)=t\log\bigl{(}f^{\#}(z)\bigr{)}$ (see (5.1) for the definition of $f^{\#}(z)$). With the above conventions, we state the following theorem.
Theorem C.
There exists an algorithm which, on input oracles of the coefficients of a hyperbolic rational map $f$ of degree $d\geqslant 2$ and an oracle of $t\in\mathbb{R}$, computes the unique equilibrium state of $f|_{\mathcal{J}_{f}}$ for the potential $t\varphi_{f}(z)$. Moreover, the Hausdorff dimension of the Julia set $\mathcal{J}_{f}$ is computable.
Note that the proof of Theorem C depends on the computability of the hyperbolic Julia sets ([Br04]). Moreover, since the computability of the topological pressure is a consequence of the proof of Theorem 1.1, the pressure function $\mathrm{P}(t)\coloneqq P(f|_{\mathcal{J}_{f}},-t\varphi_{f})$ is computable. Because $\mathrm{P}(t)$ is strictly decreasing with respect to $t\in\mathbb{R}$, as a byproduct, we can compute its unique zero, that is, by Bowen’s formula (see for example, [PU10, Theorem 9.1.6 and Corollary 9.1.7]), the Hausdorff dimension of $\mathcal{J}_{f}$. Non-uniqueness of equilibrium states. Determining which dynamical system admits a unique equilibrium state is a central problem in ergodic theory and a valuable tool for studying the statistical properties of a system (for example, detecting the existence of a phase transition). As indicated by Hofbauer [Ho75], this question was motivated by the studies of intrinsically ergodic dynamical systems, i.e., those having a unique measure of maximal entropy, see for examples, [Bu97, Bu05, Bo74, CT12, Ho79, Ho81, Pa64, Weis70, Weis73]. For general equilibrium states beyond measures of maximal entropy, most results focus on proving the uniqueness, and there are limited examples of systems and potentials with multiple equilibrium states. In the context of a uniformly hyperbolic (or expanding) map, Bowen [Bo75] proved that an equilibrium state does exist and is unique if the potential is Hölder continuous and the map is topologically transitive. In addition, the theory for finite shifts was developed and used to achieve similar results for smooth dynamics. In contrast, Hofbauer [Ho75] constructed a particular class of continuous but non-Hölder potentials $\phi$ for the full shift, each of which admits two equilibrium states. Beyond uniform hyperbolicity, the uniqueness of the equilibrium state was studied by many authors, including Bruin, Keller, Li, Rivera-Letelier, Iommi, Doobs, and Todd [BK98, LRL14, IT15, IT10, DT21] for interval maps; Denker and Urbanski [DU91] for rational maps; the third-named author, Das, Przytycki, Tiozzo, Urbański, and Zdunik [Li18, DPTUZ21] for branched covering maps; Buzzi, Sarig, and Yuri [BS03, Yu03], for countable Markov shifts and for piecewise expanding maps in one and higher dimensions. For local diffeomorphisms with some non-uniform expansion, there are results due to Varandas and Viana [VV10], Pinheiro [Pi11], Ramos and Viana [RV17]. All of these results focus on establishing the existence and uniqueness of the equilibrium state for a potential with low oscillation. On the other hand, much fewer examples that admit multiple equilibrium states have been explored. They include
  1. (i)
    Makarov and Smirnov [MS00, MS03], and Rivera-Letelier and Przytycki [PRL11] systematically studied the pressure function for rational maps $f$ with geometric potential $\phi_{t}=-t\log\lvert f^{\#}\rvert$, and constructed examples admitting multiple equilibrium states. For a similar study on (generalized) interval maps, see Rivera-Letelier and Przytycki [PRL14].
  2. (ii)
    Examples of intermittent maps admitting multiple equilibrium states for the geometric potential were studied in [VV10] via Pesin theory.
Applying Computable Analysis, we establish the following result.
Theorem D.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $X$ a recursively compact set. Assume that $T\colon X\rightarrow X$ is a computable map with $h_{\operatorname{top}}(T)=0$ satisfying that there exists no computable point in the non-wandering set $\Omega(T)$. Let $\varphi\colon X\rightarrow\mathbb{R}$ be a computable function satisfying that the pressure $P(T,\varphi)$ is computable. Then there exist at least two equilibrium states of $T$ for $\varphi$.
We refer to (2.1) for the definition of non-wandering sets. For example, it follows immediately from Theorem D and the construction in [GHR10, Subsection 4.1] that given a number $\epsilon>0$, there exists a computable function $T\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ from [GHR10, Subsection 4.1] satisfying that for each arc $I$ larger than $\epsilon$ with computable endpoints, there exist at least two equilibrium states of $T$ for the potential $-d_{I}$, where $d_{I}(x)$ denotes the Euclidean distance between $x$ and $I$ along the circle $\mathbb{S}^{1}$ for each $x\in\mathbb{S}^{1}$. More generally, we have the following result.
Theorem E.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $X$ a recursively compact set. Assume that $T\colon X\rightarrow X$ is a computable map satisfying that the measure-theoretic entropy function $\mathcal{H}\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ given by
$\mathcal{H}(\mu)\coloneqq\begin{cases}h_{\mu}(T)&\text{if }\mu\in\mathcal{M}_{T}(X);\\ 0&\text{if }\mu\notin\mathcal{M}_{T}(X)\end{cases}$
is upper-computable on $\mathcal{M}_{T}(X)$, and $\varphi\colon X\rightarrow\mathbb{R}$ is a computable function satisfying that the topological pressure $P(T,\varphi)$ is lower-computable. Then the set consisting of all equilibrium states of $T$ for $\varphi$ is recursively compact. Consequently, if there exists a non-computable equilibrium state of $T$ for $\varphi$, then there exists at least two equilibrium states of $T$ for $\varphi$.
Here $\mathcal{M}_{T}(X)$ is the space of $T$-invariant Borel probability measures equipped with the Wasserstein–Kantorovich metric $W_{\rho}$ (see Section 2.1). Counterexamples. In the following theorem, we provide an example of a computable dynamical system with zero topological entropy (hence, for which equilibrium states exist), whose every equilibrium state for every continuous potential is non-computable.
Theorem F.
There exists a computable homeomorphism $T\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ of the unit circle $\mathbb{S}^{1}$ such that for each continuous potential $\varphi\colon\mathbb{S}^{1}\rightarrow\mathbb{R}$, the following statements hold:
  1. (i)
    There exists at least one equilibrium state of $T$ for $\varphi$.
  2. (ii)
    There is no computable $T$-invariant probability measure.
Recall that the topological entropy reflects a degree of complexity of a dynamical system. We, furthermore, show that there is a dynamical system on the $2$-torus with arbitrarily high topological entropy that admits equilibrium states for each Hölder continuous potential, and whose equilibrium states are all non-computable.
Theorem G.
There exists a continuous map $\widehat{T}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}\times\mathbb{S}^{1}$ with non-zero topological entropy such that for each continuous potential $\widehat{\varphi}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{R}$, the following statements hold:
  1. (i)
    There exists at least one equilibrium state of $\widehat{T}$ for $\widehat{\varphi}$.
  2. (ii)
    There is no computable $\widehat{T}$-invariant probability measure.
Strategy of the proof. We discuss below the strategies applied in the proofs of our main results. Let us first concentrate on the proof of Theorem A. Here is a precise version of Theorem A.
Theorem 1.1.
There exists an algorithm that satisfies the following property: For each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, this algorithm outputs a Borel probability measure $\mu_{n}\in\mathcal{P}(X)$ supported on finitely many points of $\mathcal{S}$ such that the Wasserstein–Kantorovich distance
$W_{\rho}(\mu_{n},\mu)\leqslant 2^{-n},$
after inputting the following data in this algorithm:
  1. (i)
    an algorithm outputting a net $G\subseteq\mathcal{S}$ of $X$ with any given precision (in the sense of Definition 3.16),
  2. (ii)
    two algorithms computing functions $\varphi$ and $T$ (in the sense of Definition 3.9), respectively,
  3. (iii)
    two rational constants $a_{0}$ and $v_{0}$ satisfying that $\varphi$ is Hölder continuous with exponent $v_{0}$ and constant $a_{0}$ (in the sense of Definition 2.6),
  4. (iv)
    three rational constants $\eta$, $\lambda,$ and $\xi$ satisfying that $T$ is distance-expanding with these constants (in the sense of Definition 2.1),
  5. (v)
    a constant $n\in\mathbb{N}$.
The Wasserstein–Kantorovich metric $W_{\rho}$ on the set $\mathcal{P}(X)$ of Borel probability measures is defined in Section 2.1. Our proof of Theorem 1.1 is inspired by the proof of Theorem A in [BBRY11]. However, since we work on equilibrium states, additional methods are required. The most crucial of them is the technique of cones. We use it in Subsection 4.1 to estimate the convergence rate of the iterations of a normalized Ruelle–Perron–Frobenius operator (see the definition in Subsection 2.3) with explicit expressions for related constants (see Theorem 4.1). In the setting of smooth (or more generally Markov) expanding maps with smooth observable, related bounds were investigated by Ferrero ([Fe81]), Rychlik ([Ry89]), and Hunt ([Hu96]). In two groundbreaking papers [Liv95a, Liv95b] in 1995, Liverani used the technique of cones (based on an earlier work of Birkhoff [Bi57]) and gave a new upper bound of the convergence rate of the iterations of a normalized Ruelle–Perron–Frobenius operator for general (non-Markov) non-smooth hyperbolic systems with discontinuities. A critical ingredient of this method is the existence of a convex cone of functions that is mapped strictly inside itself by the Ruelle–Perron–Frobenius operator. Moreover, the diameter of the image with respect to some projective metric gives an estimate for the upper bound of the convergence rate (see Proposition 4.3). In the setting of smooth expanding maps on compact connected manifolds, Viana [Vi97, Chapter 2] provides a comprehensive estimate for the convergence rate of the iterations of a normalized Ruelle–Perron–Frobenius operator in Hölder norm. Notably, because of the connectedness, some crucial local estimates of the oscillations of the functions in the cone can be extended globally (to the whole manifold). This, however, cannot work in our setting, since we allow for general disconnected spaces. To remedy this, we add a new parameter into the definition of cones used in [Vi97], effectively restricting to smaller cones of functions (see Definition 4.4), in order to estimate the diameters (see Proposition 4.6). This extra parameter takes care of the required global bounds on oscillations. Moreover, to make sure that the operator is a strict contraction on these new cones, we use a different version of the normalized Ruelle–Perron–Frobenius operator and utilize additional techniques (see Proposition 4.5). Finally, we use the convergence rate of the new operator to obtain the desired bounds for the normalized Ruelle–Perron–Frobenius operator. We need this new finer estimate for the convergence rate (Theorem 4.1) to establish the computability of the eigenfunction $u_{\varphi}$ of the Ruelle–Perron–Frobenius operator. Furthermore, in Subsection 4.2, we deduce the computability of the Jacobian of the equilibrium state whose explicit expression is given by
$J_{\varphi}(x)\coloneqq\frac{u_{\varphi}(T(x))}{u_{\varphi}(x)}\exp(P(T,\varphi)-\varphi(x)).$
In Subsection 4.3, we establish the lower-computable openness of some subsets $U$ and $W$ of $\mathcal{P}(X)$ by expressing them as unions of uniformly computable sequences of lower-computable open sets. It is worth noting that the lower-computable openness of $U$ depends on the computability of $T$, while the lower-computable openness of $W$ depends on the computability of the Jacobian function $J_{\varphi}(x)$ in addition to the computability of $T$. In Subsection 4.4, we complete the proof of Theorem 1.1 and apply it to establish Theorem B. Here is a precise version of Theorem B.
Theorem 1.2.
There exists an algorithm that satisfies the following property: For each septet $(X,\,D,\,\rho,\,\mathcal{S},\,\mathcal{S}^{\prime},\,\varphi,\,T)$ satisfying the Assumption B in Section 1, this algorithm outputs $\mu_{n}\in\mathcal{P}(X)$ supported on finitely many points of $\mathcal{S}$ such that the Wasserstein–Kantorovich distance
$W_{\rho}(\mu_{n},\mu)\leqslant 2^{-n},$
after inputting the following data in this algorithm:
  1. (i)
    an algorithm outputting a net $G\subseteq\mathcal{S}^{\prime}$ of $D$ with any given precision (in the sense of Definition 3.16),
  2. (ii)
    two algorithms computing functions $\varphi$ and $T|_{D}$ (in the sense of Definition 3.9), respectively,
  3. (iii)
    two rational constants $a_{0}$ and $v_{0}$ satisfying that $\varphi$ is Hölder continuous with exponent $v_{0}$ and constant $a_{0}$ (in the sense of Definition 2.6),
  4. (iv)
    three rational constants $\eta$, $\lambda,$ and $\xi$ satisfying that $T$ is distance-expanding with these constants (in the sense of Definition 2.1),
  5. (v)
    a constant $n\in\mathbb{N}$.
Next, in Section 5, we discuss the strategy of our proof of Theorem C. By Theorem 1.2, it suffices to check that the septet (^$\bigl{(}\widehat{\mathbb{C}},\,\mathcal{J}_{f},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2},\,\bigcup_{n\in\mathbb{N}}f^{-n}(w),\,t\varphi_{f},\,f\bigr{)}$ satisfies the Assumption B in Section 1, where $\mathbb{Q}^{2}$ is defined as $\{a+bi:a,\,b\in\mathbb{Q}\}$, and $w$ is a repelling periodic point of the rational map $f$. The main difficulty here is to show that $\mathcal{J}_{f}$ is recursively compact in the computable metric space $\bigl{(}\mathcal{J}_{f},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}f^{-n}(w)\bigr{)}$. We use the computability of the hyperbolic Julia sets proved by Braverman [Br04] to establish this statement. All of our analyses for Theorems 1.1 and 1.2 should still work after necessary modifications if we replace the algorithms in the inputs by oracles. For the simplicity of the presentation, we leave the details to the reader. Structure of the paper. In Section 2, we introduce our notations and review some basic notions and results in ergodic theory and thermodynamic formalism in particular. Section 3 covers some basics of Computable Analysis. Section 4 is dedicated to the proofs of Theorems 1.1 and 1.2. In Subsection 4.1, we prove that the convergence of the iterations of a normalized Ruelle–Perron–Frobenius operator is at an exponential rate with explicit expressions for the related constants. Based on this, we demonstrate the computability of topological pressure $P(T,\varphi)$ and the function $u_{\varphi}$, which leads to the computability of the Jacobian of the equilibrium state in Subsection 4.2. Then, we prove Lemma 4.13 in Subsection 4.3 and establish Theorems 1.1 and 1.2 in Subsection 4.4. In Section 5, we introduce some results on the dynamical and algorithmic aspects of hyperbolic rational maps and prove Theorem C. In Section 6, we establish Theorem E, from which we derive Theorem D. In Section 7, we investigate and construct some computable systems whose equilibrium states are all non-computable, establishing Theorems F and G. In Subsection 7.1, we recall that the computable map $T\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ given in [GHR10, Subsection 4.1] satisfies statements in Theorem F. Finally, in Subsection 7.2, we construct $\widehat{T}\coloneqq T\times T_{d}$ to establish Theorem G, where $T_{d}(x)\coloneqq dx\pmod{1}$ for each $d\in\mathbb{N}$ and each $x\in\mathbb{S}^{1}$.
Acknowledgments
I. Binder was partially supported by an NSERC Discovery grant. Z. Li and Q. He were partially supported by NSFC Nos. 12101017, 12090010, 12090015, and BJNSF No. 1214021. Q. He was also supported by Peking University Funding Nos. 7101303303 and 6201001846. Y. Zhang was partially supported by NSFC Nos. 12161141002, 11871262, and 12271432.
2. Preliminaries on thermodynamic formalism
In this section, we go over some notations, key concepts, and useful results in ergodic theory and thermodynamic formalism in particular.
2.1. Notations
Denote {1, 2, $\mathbb{N}\coloneqq\{1,\,2,\,3,\,\ldots\}$ and $\mathbb{N}_{0}\coloneqq\{0\}\cup\mathbb{N}$. For each $k\in\mathbb{N}$, we fixed a bijection between $\mathbb{N}^{k}$ and $\mathbb{N}$ call it the “canonical bijection between $\mathbb{N}^{k}$ and $\mathbb{N}$”. Set *$\mathbb{N}^{*}\coloneqq\bigcup_{k\in\mathbb{N}}\mathbb{N}^{k}$, and fix a bijection between $\mathbb{N}^{*}$ and $\mathbb{N}$ that we call it the “canonical bijection between $\mathbb{N}^{*}$ and $\mathbb{N}$”. For a subset $A\subseteq X$, the characteristic function of $A$ is denoted by $\mathbbm{1}_{A}$. We write $\mathbbm{1}\coloneqq\mathbbm{1}_{X}$. Let $(X,\rho)$ be a compact metric space. Let $C(X)$ and $C^{*}(X)$ denote the Banach space of continuous functions from $X$ to $\mathbb{R}$ equipped with the norm $\|u\|_{\infty}\coloneqq\sup\{\lvert u(x)\rvert:x\in X\}$ and the dual space of $C(X)$, respectively. Given a constant $c>0$, a function $g\in C(X)$ is $c$-Lipschitz if $\lvert g(x)-g(y)\rvert\leqslant c\rho(x,y)$ for each pair of $x,\,y\in X$. We use $c\operatorname{-Lip}(X)$ to denote the set of real-valued $c$-Lipschitz continuous functions, and $B(x,r)=B_{\rho}(x,r)$ to denote the open ball of radius $r>0$ centered at $x\in X$.
Let $T\colon X\rightarrow X$ be a continuous map. The non-wandering set for $T$ is denoted by
(2.1) $\Omega(T)\coloneqq\Bigl{\{}x\in X:\bigcup_{n\in\mathbb{N}}T^{-n}(B(x,r))\cap B(x,r)\neq\emptyset\text{ for each }r>0\Bigr{\}}.$
For all $\varphi\in C(X)$, $n\in\mathbb{N}$, and $x\in X$, denote
$S_{n}\varphi(x)\coloneqq\sum\limits_{k=0}^{n-1}\varphi\circ T^{k}(x).$
Let $\mathcal{P}(X)$ and $\mathcal{M}_{T}(X)$ denote the set of Borel probability measures and the set of $T$-invariant Borel probability measures on $X$, respectively. The Wasserstein–Kantorovich metric $W_{\rho}$ on $\mathcal{P}(X)$ is defined by
(2.2) $W_{\rho}(\mu,\nu)\coloneqq\sup\biggl{\{}\biggl{\lvert}\int\!f\,\mathrm{d}\mu-\int\!f\,\mathrm{d}\nu\biggr{\rvert}:f\in 1\operatorname{-Lip}(X)\biggr{\}}.$
2.2. Basic concepts in ergodic theory
We begin with a brief introduction of distance-expanding maps and topologically transitive (resp. exact) maps. See for example, [PU10, Chapters 3 and 4] for more details. Given constants $\eta>0$ and $\lambda>1$, assume that $(X,\rho)$ is a compact metric space and $T\colon X\rightarrow X$ is a continuous open map satisfying
(2.3) $\rho(T(x),T(y))\geqslant\lambda\rho(x,y)\quad\text{for each pair of }x,\,y\in X\text{ with }\rho(x,y)\leqslant 2\eta.$
Then there exists a constant $\xi>0$ such that $B(T(x),\xi)\subseteq T(B(x,\eta))$ for each $x\in X$. Hence, for each $x\in X$ and each $n\in\mathbb{N}$, one can define the branches of the inverse map $T_{x}^{-n}\colon B(T^{n}(x),\xi)\rightarrow B(x,\eta)$ of $T^{n}$ in the sense of
$T^{-n}(A)=\bigcup_{x\in T^{-n}(y)}T^{-n}_{x}(A)$
for all $y\in X$ and $A\subseteq B(y,\xi)$ such that the following inequalities hold (see for example, [PU10, Section 4.1]):
(2.4) $\rho(T_{x}^{-n}(y),T_{x}^{-n}(z))\leqslant\lambda^{-n}\rho(y,z)\text{ for all }n\in\mathbb{N},\,x\in X,\,\text{and }y,\,z\in B(T^{n}(x),\xi).$
Definition 2.1.
Let $(X,\rho)$ be a compact metric space. Consider $\eta>0$, $\lambda>1$, and $\xi>0$. A map $T\colon X\rightarrow X$ is said to be distance-expanding with respect to the metric $\rho$ with constants $\eta$, $\lambda$, and $\xi$ if (2.3) and (2.4) both hold.
In a topological space $X$, recall that a continuous map $T\colon X\rightarrow X$ is topologically transitive if, for each pair of non-empty open sets $U,\,V\subseteq X$, there exists $n\in\mathbb{N}$ such that $T^{n}(U)\cap V\neq\emptyset$. Moreover, as a stronger property, we say a continuous map $T\colon X\rightarrow X$ is topologically exact if, for each non-empty open sets $U\subseteq X$, there exists $n\in\mathbb{N}$ such that $T^{n}(U)=X$. Under the above conventions, given constants $\eta>0$, $\lambda>1$, and $\xi>0$, a list of assumptions that are applied in the rest of this section is displayed below:
Assumption C.
  1. (i)
    The metric space $(X,\rho)$ is compact.
  2. (ii)
    The map $T\colon X\rightarrow X$ is continuous, topologically transitive, open, and distance-expanding with respect to the metric $\rho$ constants $\eta$, $\lambda$, and $\xi$.
Next, we recall some notations and results in ergodic theory. For more details, we refer to [PU10, Chapter 2]. Let $X$ be a topological space. A measurable partition $\mathcal{A}$ of $X$ is a cover $\mathcal{A}=\{A_{j}:j\in J\}$ of $X$ consisting of finite or countably many mutually disjoint Borel sets. Unless otherwise stated, we assume that our partitions are all measurable. For each partition $\mathcal{A}$ of $X$ and $x\in X$, the unique element of $\mathcal{A}$ containing $x$ is denoted by $\mathcal{A}(x)$. Let $\mathcal{A}=\{A_{j}:j\in J\}$ and =$\mathcal{B}=\{B_{k}:k\in K\}$ be two covers of $X$, respectively, where $J$ and $K$ are corresponding index sets. We say that $\mathcal{A}$ is a refinement of $\mathcal{B}$ if for each $j\in J$, there exists $k=k(j)\in K$ such that $A_{j}\subseteq B_{k}$. The common refinement $\mathcal{A}\vee\mathcal{B}$ of $\mathcal{A}$ and $\mathcal{B}$ is a cover given by $\mathcal{A}\vee\mathcal{B}\coloneqq\{A_{j}\cap B_{k}:j\in J,\,k\in K\}.$ For each continuous map $T\colon X\rightarrow X$, by $T^{-1}(\mathcal{A})$ we denote the partition $\{T^{-1}(A_{j}):j\in J\}$, and for each $n\in\mathbb{N}$, write
$\mathcal{A}_{T}^{n}\coloneqq\bigvee\limits_{j=0}^{n-1}T^{-j}(\mathcal{A})=\mathcal{A}\vee T^{-1}(\mathcal{A})\vee\cdots\vee T^{-(n-1)}(\mathcal{A}).$
Definition 2.2.
Let $(X,\rho)$ be a compact metric space, $T\colon X\rightarrow X$ be a continuous map, and $\varphi\colon X\rightarrow\mathbb{R}$ be a continuous function (which is often called a potential). Denote by $\mathcal{Y}$ the set of the sequences $\{\mathcal{V}_{n}\}_{n\in\mathbb{N}}$ of open finite covers of $X$ satisfying that the meshes of the covers $\mathcal{V}_{n}$ converge to zero as $n\rightarrow+\infty$, that is,
$\mathcal{Y}\coloneqq\bigl{\{}\{\mathcal{V}_{n}\}_{n\in\mathbb{N}}:\lim\limits_{n\to+\infty}\operatorname{mesh}(\mathcal{V}_{n})=0\bigr{\}},$
where the mesh of a cover $\mathcal{V}$ of $X$ is given by
$\operatorname{mesh}(\mathcal{V})\coloneqq\sup\{\operatorname{diam}(V):V\in\mathcal{V}\}.$
Define the topological pressure as the supremum of the following limits taken over all sequences $\{\mathcal{V}_{n}\}\in\mathcal{Y}$:
(2.5) $\lim\limits_{n\to+\infty}\lim\limits_{m\to+\infty}\frac{1}{n}\log\biggl{(}\inf_{\mathcal{V}}\biggl{\{}\sum_{U\in\mathcal{V}}\exp(\sup\{S_{m}\varphi(x):x\in U\})\biggr{\}}\biggr{)},$
where $\mathcal{V}$ ranges over all covers of $X$ contained (in the sense of inclusion) in $(\mathcal{V}_{n})_{T}^{m}$. In particular, if the potential $\varphi$ is identically zero, then the topological pressure $P(T,\varphi)$ is also called the topological entropy of $T$, and is denoted by $h_{\operatorname{top}}(T)$.
We refer to [PU10, Subsection 3.2] to demonstrate that the topological pressure is well-defined by (2.5). Let $(X,\rho)$ be a compact metric space. For all continuous map $T$ of $X$ and $T$-invariant Borel probability measure $\mu\in\mathcal{M}_{T}(X)$, one can define measure-theoretic entropy $h_{\mu}(T)$ of $T$ with respect to $\mu$ (see [PU10, Chapter 2.4]). In particular, assume that $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2, and $\mu$ is a $T$-invariant Borel probability measure. Recall that a measurable function $J\colon X\rightarrow[0,+\infty)$ is a Jacobian of $T$ with respect to $\mu$ if for each $\mu$-measurable set $A\subseteq X$ on which $T$ is injective, $T(A)$ is still $\mu$-measurable and $\mu(T(A))=\int_{A}\!J\,\mathrm{d}\mu.$ By [PU10, Proposition 2.9.5], there exists a unique Jacobian $J_{\mu}$ of $T$ with respect to $\mu$. Then the measure-theoretic entropy of $T$ with respect to $\mu$ has the following explicit expression $h_{\mu}(T)=\int\!\log(J_{\mu})\,\mathrm{d}\mu$ (see for example, [PU10, Proposition 2.9.7]). Under the above notations, one has the following theorem (see for example, [PU10, Theorem 3.4.1]).
Theorem 2.3 (Variational Principle).
Let $(X,\rho)$ be a compact metric space. Then for each continuous map $T\colon X\rightarrow X$ and each potential $\varphi\in C(X)$, we have
$P(T,\varphi)=\sup\left\{h_{\mu}(T)+\int\!\varphi\,\mathrm{d}\mu:\mu\in\mathcal{M}_{T}(X)\right\}.$
In particular, if $\varphi\equiv 0$, then $h_{\operatorname{top}}(T)=\sup\{h_{\mu}(T):\mu\in\mathcal{M}_{T}(X)\}$.
Definition 2.4.
Let $(X,\rho)$ be a compact metric space. Then for each continuous map $T\colon X\rightarrow X$ and each potential $\varphi\in C(X)$, we say that a measure $\mu\in\mathcal{M}_{T}(X)$ is an equilibrium state of the map $T$ for the potential $\varphi$ if
$P(T,\varphi)=h_{\mu}(T)+\int\!\varphi\,\mathrm{d}\mu.$
In particular, for $\varphi\equiv 0$, an equilibrium state of $T$ for $\varphi$ is also called a measure of maximal entropy of $T$.
2.3. Ruelle–Perron–Frobenius operators and Gibbs states
In this subsection, we review some notions and results for the Ruelle–Perron–Frobenius operator, the Gibbs state, and the equilibrium state. Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2, and the potential $\varphi\colon X\rightarrow\mathbb{R}$ is a continuous function. Recall that the Ruelle operator $\mathcal{L}_{\varphi}$ acting on $C(X)$ is given by
(2.6) $\mathcal{L}_{\varphi}(u)(x)\coloneqq\sum\limits_{y\in T^{-1}(x)}u(y)\exp(\varphi(y))$
for each $u\in C(X)$ and each $x\in X$. By [PU10, Theorem 5.2.8], there exists an eigenmeasure $m\in\mathcal{P}(X)$ and an eigenvalue $c>0$ of the adjoint operator $\mathcal{L}_{\varphi}^{*}$ of $\mathcal{L}_{\varphi}$, i.e., $\mathcal{L}^{*}_{\varphi}(m)=cm$. Moreover, by [PU10, Proposition 5.2.11], the constant $c$ is precisely $\exp(P(T,\varphi)).$ It is convenient to consider the normalized Ruelle–Perron–Frobenius operator $\mathcal{L}_{\overline{\varphi}}$ with $\overline{\varphi}\coloneqq\varphi-P(T,\varphi).$ By (2.6), we have
(2.7) $\mathcal{L}_{\overline{\varphi}}^{*}(m)=e^{-P(T,\varphi)}\mathcal{L}_{\varphi}^{*}(m)=m.$
A measure $\mu\in\mathcal{P}(X)$ is a Gibbs state for $T$ and $\varphi$ if there exist two constants $P\in\mathbb{R}$ and $C\geqslant 1$ such that
(2.8) $C^{-1}\leqslant\frac{\mu(T^{-n}_{x}(B(T^{n}(x),\xi)))}{\exp(S_{n}\varphi(x)-Pn)}\leqslant C$
for all $x\in X$ and $n\in\mathbb{N}.$ The following result (see for example, [PU10, Propositions 4.4.3, 5.1.5, 5.2.10, Corollaries 5.2.13, Lemma 5.6.1, and Theorem 5.6.2]) implies that a $T$-invariant Gibbs state is an ergodic equilibrium state of $T$ for the potential $\varphi$.
Proposition 2.5.
Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2. Then for all $\mu\in\mathcal{M}_{T}(X)$ and Hölder continuous potential $\varphi\colon X\rightarrow\mathbb{R}$, the following statements are equivalent:
  1. (i)
    $\mu$ is a Gibbs state for $T$ and $\varphi$.
  2. (ii)
    $\mu$ is an ergodic equilibrium state of $T$ for $\varphi$.
  3. (iii)
    $\exp(-\psi)$ is a Jacobian of $T$ with respect to $\mu$, where
    $\psi\coloneqq\varphi-P(T,\varphi)-\log(u_{\varphi}\circ T)+\log(u_{\varphi}).$
    Here $u_{\varphi}$ is a subsequential limit of the sequence $\bigl{\{}\frac{1}{n}\sum_{j=0}^{n-1}\mathcal{L}_{\overline{\varphi}}^{j}(\mathbbm{1})\bigr{\}}_{n\in\mathbb{N}}$.
The existence and uniqueness of the subsequential limit $u_{\varphi}$ is well-known. We include a proof in Corollary 2.9 together with some quantitative bounds specialized for the purpose of Computable Analysis. In the remainder of this subsection, we quantitatively re-develop some familiar facts about the Ruelle–Perron–Frobenius operators from the perspective of Computable Analysis in preparation for more technical analysis in Section 4.
Definition 2.6.
Let $(X,\rho)$ be a compact metric space, and $v_{0}\in(0,1]$ and $a_{0}>0$ be constants. A function $\varphi\colon X\rightarrow\mathbb{R}$ is Hölder continuous (with respect to the metric $\rho$) with exponent $v_{0}$ and constant $a_{0}$ if $\lvert\varphi(x)-\varphi(y)\rvert\leqslant a_{0}\rho(x,y)^{v_{0}}$ for all $x,\,y\in X$.
Lemma 2.7.
Fix arbitrary constants $\eta,\,\xi,\,a_{0},\,v_{0}>0$, and $\lambda>1$. Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2 with constants $\eta$, $\lambda$, and $\xi$, and the potential $\varphi\colon X\rightarrow\mathbb{R}$ be Hölder continuous with respect to the metric $\rho$ with exponent $v_{0}$ and constant $a_{0}$. Define $\overline{\varphi}(x)\coloneqq\varphi(x)-P(T,\varphi)$ for each $x\in X$, and $a\coloneqq\frac{a_{0}}{\lambda^{v_{0}}-1}$. Then
(2.9) $\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)\leqslant\exp(a\rho(x,y)^{v_{0}})\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y)$
for each $n\in\mathbb{N}$ and each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$.
Proof.
Since $T$ is an open map that is distance-expanding with constants $\eta$, $\lambda$, and $\xi$. Then for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$ and each $\overline{x}\in T^{-n}(x)$, one can find a point $y(\overline{x})\coloneqq T_{\overline{x}}^{-n}(y)\in T^{-n}(y)$ satisfying that $\rho(T^{i}(\overline{x}),T^{i}(y(\overline{x})))\leqslant\lambda^{i-n}\rho(x,y)\leqslant\xi$ for each integer $10\leqslant i\leqslant n-$ 1. Noting that $\overline{\varphi}(x)=\varphi(x)-P(T,\varphi)$ is Hölder continuous with exponent $v_{0}$ and constant $a_{0}$, we have
$S_{n}\overline{\varphi}(\overline{x})-S_{n}\overline{\varphi}(y(\overline{x}))$ $=\sum_{i=0}^{n-1}(\overline{\varphi}(T^{i}(\overline{x}))-\overline{\varphi}(T^{i}(y(\overline{x}))))$
$\leqslant\sum_{i=0}^{n-1}a_{0}\rho(T^{i}(\overline{x}),T^{i}(y(\overline{x})))^{v_{0}}$
$\leqslant a_{0}\rho(x,y)^{v_{0}}\cdot\sum_{i=0}^{n-1}\lambda^{(i-n)v_{0}}$
$\leqslant a_{0}\rho(x,y)^{v_{0}}/(\lambda^{v_{0}}-1)$
for each $\overline{x}\in T^{-n}(x)$. Together with (2.6), we can conclude that
$\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)$ $=\sum_{\overline{x}\in T^{-n}(x)}\exp(S_{n}\overline{\varphi}(\overline{x}))$
$\leqslant\sum_{\overline{x}\in T^{-n}(x)}\exp(S_{n}\overline{\varphi}(y(\overline{x})))\cdot\exp(a_{0}\rho(x,y)^{v_{0}}/(\lambda^{v_{0}}-1))$
$\leqslant\exp(a_{0}\rho(x,y)^{v_{0}}/(\lambda^{v_{0}}-1))\mathcal{L}^{n}_{\overline{\varphi}}(\mathbbm{1})(y)$
for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$. ∎
Lemma 2.8.
Fix arbitrary constants $\eta,\,\xi,\,a_{0},\,v_{0}>0$, and $\lambda>1$. Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2 with constants $\eta$, $\lambda$, and $\xi$, and the potential $\varphi\colon X\rightarrow\mathbb{R}$ be Hölder continuous with respect to the metric $\rho$ with exponent $v_{0}$ and constant $a_{0}$. Define $\overline{\varphi}(x)\coloneqq\varphi(x)-P(T,\varphi)$ for each $x\in X$, $a\coloneqq\frac{a_{0}}{\lambda^{v_{0}}-1}$, and $D\coloneqq\max_{x\in X}\operatorname{card}\bigl{(}T^{-1}(x)\bigr{)}$. Denote by $G=\{x_{1},\,x_{2},\,\ldots,\,x_{l}\}$ a $\xi$-net of $X$. Then for all $x,\,y\in X$, and $n\in\mathbb{N}$, we have
(2.10) $\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)\leqslant C\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y)$
and
(2.11) $C^{-1}\leqslant\mathcal{L}^{n}_{\overline{\varphi}}(\mathbbm{1})(x)\leqslant C.$
Here $C\coloneqq D^{N}\exp(4a\xi^{v_{0}}+2N\|\varphi\|_{\infty})>1$, where $N\in\mathbb{N}$ is a constant satisfying that
(2.12) $\bigcup_{k=0}^{N}T^{k}(B(x_{i},\xi))=X$
for each integer $i\in[1,l]$.
See [PU10, Theorem 4.3.12 (2)] for the existence of $N$.
Proof.
By Lemma 2.7, we have
(2.13) $\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)\big{/}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y)\leqslant\exp(a\rho(x,y)^{v_{0}})\leqslant\exp(a\xi^{v_{0}}).$
for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$ and each $n\in\mathbb{N}$. Now we are ready to demonstrate (2.10). Let $x,\,y\in X$. Since $G$ is a $\xi$-net, there exist two integers $i,\,j\in[1,l]$ satisfying that $\rho(x_{i},x)<\xi$ and $\rho(x_{j},y)<\xi$. Hence, by (2.12), there exists an integer $m\in[0,N]$ such that $T^{m}(B(x_{i},\xi))\cap B(x_{j},\xi)\neq\emptyset$. Now let $u$ be a point in $T^{-m}(B(x_{j},\xi))\cap B(x_{i},\xi)$. By (2.6) and (2.13), then for each $n\in\mathbb{N}$, we have
$\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x_{i})$ $\leqslant\exp(a\xi^{v_{0}})\cdot\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(u)$
$=\exp(a\xi^{v_{0}})\sum_{\overline{u}\in T^{-n}(u)}\exp(S_{n+m}\overline{\varphi}(\overline{u})-S_{m}\overline{\varphi}(u))$
$\leqslant\exp\bigl{(}a\xi^{v_{0}}-m\inf_{x\in X}\overline{\varphi}(x)\bigr{)}\sum_{\overline{u}\in T^{-(n+m)}(T^{m}(u))}\exp(S_{n+m}\overline{\varphi}(\overline{u}))$
$=\exp\bigl{(}a\xi^{v_{0}}-m\inf_{x\in X}\overline{\varphi}(x)\bigr{)}\sum_{\overline{u}\in T^{-(n+m)}(T^{m}(u))}\exp(S_{n}\overline{\varphi}(T^{m}(\overline{u}))+S_{m}\overline{\varphi}(\overline{u}))$
$\leqslant\exp\bigl{(}a\xi^{v_{0}}+m\sup_{x\in X}\overline{\varphi}(x)-m\inf_{x\in X}\overline{\varphi}(x)\bigr{)}\sum_{\overline{u}\in T^{-(n+m)}(T^{m}(u))}\exp(S_{n}\overline{\varphi}(T^{m}(\overline{u})))$
$\leqslant\exp(a\xi^{v_{0}}+2N\|\overline{\varphi}\|_{\infty})D^{N}\sum_{\overline{u}\in T^{-n}(T^{m}(u))}\exp(S_{n}\overline{\varphi}(\overline{u}))$
$=\exp(a\xi^{v_{0}}+2N\|\varphi\|_{\infty})D^{N}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(T^{m}(u))$
$\leqslant\exp(2a\xi^{v_{0}}+2N\|\varphi\|_{\infty})D^{N}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x_{j}).$
Therefore, by (2.13), it follows from $\rho(x_{i},x)<\xi$ and $\rho(x_{j},y)<\xi$ that for each $n\in\mathbb{N}$,
$\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)$ $\leqslant\exp(a\xi^{v_{0}})\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x_{i})$
$\leqslant\exp(3a\xi^{v_{0}}+2N\|\varphi\|_{\infty})D^{N}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x_{j})$
$\leqslant\exp(4a\xi^{v_{0}}+2N\|\varphi\|_{\infty})D^{N}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y).\qed$
By (2.7), for each $n\in\mathbb{N}$,
(2.14) $\int\mathcal{L}^{n}_{\overline{\varphi}}(\mathbbm{1})\,\mathrm{d}m=\int\!\mathbbm{1}\,\mathrm{d}m=1.$
This implies that $\inf_{x\in X}\mathcal{L}^{n}_{\overline{\varphi}}(\mathbbm{1})(x)\leqslant 1\leqslant\sup_{x\in X}\mathcal{L}^{n}_{\overline{\varphi}}(\mathbbm{1})(x).$ Then by (2.10), we obtain (2.11) for all $n\in\mathbb{N}$ and $x\in X$.
Corollary 2.9.
Fix arbitrary constants $\eta,\,\xi,\,a_{0},\,v_{0}>0$, and $\lambda>1$. Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2 with constants $\eta$, $\lambda$, and $\xi$, and the potential $\varphi\colon X\rightarrow\mathbb{R}$ be Hölder continuous with respect to the metric $\rho$ with exponent $v_{0}$ and constant $a_{0}$. Define $\overline{\varphi}(x)\coloneqq\varphi(x)-P(T,\varphi)$ for each $x\in X$. Then the sequence $\bigl{\{}\frac{1}{n}\sum_{j=0}^{n-1}\mathcal{L}_{\overline{\varphi}}^{j}(\mathbbm{1})\bigr{\}}_{n\in\mathbb{N}}$ converges uniformly to a function $u_{\varphi}\in C(X)$ satisfying
(2.15) $\mathcal{L}_{\overline{\varphi}}(u_{\varphi})=u_{\varphi},$
(2.16) $u_{\varphi}(x)\leqslant\exp(a_{0}\rho(x,y)^{v_{0}}/(\lambda^{v_{0}}-1))\cdot u_{\varphi}(y)$
for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$, and
(2.17) $C^{-1}\leqslant u_{\varphi}(x)\leqslant C$
for each $x\in X$, where $C\geqslant 1$ is a constant from Lemma 2.8. Moreover, if $m\in\mathcal{P}(X)$ satisfies (2.7), then
(2.18) $\int\!u_{\varphi}\,\mathrm{d}m=1,$
and $\mu\in\mathcal{P}(X)$ given by
(2.19) $\mu(A)\coloneqq\int_{A}\!u_{\varphi}\,\mathrm{d}m,\quad\text{ for each Borel set }A\subseteq X,$
is the unique $T$-invariant Gibbs state for $T$ and $\varphi$.
Proof.
To prove this corollary, we first demonstrate (2.15), (2.16), (2.17), and (2.18) for a subsequential limit of the sequence $\bigl{\{}\frac{1}{n}\sum_{j=0}^{n-1}\mathcal{L}_{\overline{\varphi}}^{j}(\mathbbm{1})\bigr{\}}_{n\in\mathbb{N}}$, then prove that this sequence has a unique subsequential limit, and finally establish that $\mu$ is the unique $T$-invariant Gibbs state for $T$ and $\varphi$. First, by (2.9) and (2.11), we have
(2.20) $\bigl{\lvert}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)-\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y)\bigr{\rvert}\leqslant\biggl{\lvert}\frac{\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)}{\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y)}-1\biggr{\rvert}\cdot\bigl{\lvert}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y)\bigr{\rvert}\leqslant C\biggl{(}\exp\biggl{(}\frac{a_{0}\rho(x,y)^{v_{0}}}{\lambda^{v_{0}}-1}\biggr{)}-1\biggr{)}$
for all $n\in\mathbb{N}$ and $x,y\in X$ with $\rho(x,y)\leqslant\xi$. For each $n\in\mathbb{N}$, set $u_{n}\coloneqq\frac{1}{n}\sum_{j=0}^{n-1}\mathcal{L}_{\overline{\varphi}}^{j}(\mathbbm{1})$. By (2.11) and (2.20), $\{u_{n}\}_{n\in\mathbb{N}}$ is a uniformly bounded sequence of equicontinuous functions on $X$. By the Arzelà–Ascoli Theorem, there exists a continuous function $u_{\varphi}$ and an increasing sequence $\{n_{i}\}_{i\in\mathbb{N}}$ such that $u_{n_{i}}\rightarrow u_{\varphi}$ uniformly on $X$ as $i\rightarrow+\infty$. Thus it follows from the definition of $\{u_{n}\}_{n\in\mathbb{N}}$, (2.9), (2.11), and (2.14) that $u_{\varphi}$ satisfies (2.15), (2.16), (2.17), and (2.18). Next, we demonstrate that $u_{\varphi}$ is the unique subsequential limit of $\{u_{n}\}_{n\in\mathbb{N}}$. Suppose that $v_{\varphi}$ is an arbitrary subsequential limit of $\{u_{n}\}_{n\in\mathbb{N}}$. Then $v_{\varphi}$ is also a continuous function satisfying (2.15), (2.16), (2.17), and (2.18). Set
$t\coloneqq\sup\{s\in\mathbb{R}:u_{\varphi}(x)-sv_{\varphi}(x)>0\text{ for all }x\in X\}.$
Then there exists $y\in X$ such that $u_{\varphi}(y)=tv_{\varphi}(y)$. Since $\mathcal{L}^{n}_{\overline{\varphi}}(u_{\varphi}-tv_{\varphi})(y)=u_{\varphi}(y)-tv_{\varphi}(y)=0$ for each $n\in\mathbb{N}$, we obtain that $u_{\varphi}(x)-tv_{\varphi}(x)=0$ for each $x\in\bigcup_{i\in\mathbb{N}}T^{-i}(y)$. Since $T$ is topologically transitive, the set $\bigcup_{i\in\mathbb{N}}T^{-i}(y)$ is dense in $X$. Together with the continuity of $u_{\varphi}-tv_{\varphi}$, $u_{\varphi}(x)-tv_{\varphi}(x)=0$ for each $x\in X$. Because $\int\!u_{\varphi}\,\mathrm{d}m=\int\!v_{\varphi}\,\mathrm{d}m=1$, we obtain $t=1$. This implies that $u_{\varphi}=v_{\varphi}$. Therefore $u_{n}\rightarrow u_{\varphi}$ uniformly on $X$ as $n\rightarrow+\infty$. By [PU10, Proposition 5.2.11], $m$ is a Gibbs state for $T$ and $\varphi$. According to [PU10, Proposition 5.1.1] and (2.19), $\mu$ is also a Gibbs state for $T$ and $\varphi$. Recall that for each $m_{0}\in\mathcal{P}(X)$, a linear operator $\mathcal{L}_{m_{0}}\colon L^{1}(m_{0})\rightarrow L^{1}(m_{0})$ is defined in [PU10, Chapter 5.2]. Here, for the Gibbs state $m$, $\mathcal{L}_{m}$ equals exactly $\mathcal{L}_{\overline{\varphi}}$. Then by [PU10, Proposition 5.2.2], it follows from (2.15) that $\mu$ is $T$-invariant. By [PU10, Corollary 5.2.14], one can see that $\mu$ is the unique $T$-invariant Gibbs state for $T$ and $\varphi$. ∎
3. Preliminaries on Computable Analysis
In this section, we recall some notions and results in Computable Analysis. The definitions we adopted in this section are consistent with [Weih00]. Consequently, it is convenient to think of the algorithms or machines mentioned below as Type-2 machines defined in [Weih00, Definition 2.1.1]. For more details, we refer the reader to [BBRY11, Section 3], [GHR10, Section 2], and [Weih00].
3.1. Algorithms and computability over the reals
Definition 3.1.
Given $k\in\mathbb{N}$, we say that a function $f\colon\mathbb{N}^{k}\rightarrow\mathbb{Z}$ is computable, if there exists an algorithm $\mathcal{A}$ which, upon input a sequence of $k$ positive integers $\{x_{i}\}_{i=1}^{k}$, outputs the value of $f(x_{1},x_{2},\ldots,x_{k})$.
For a countable set $S$, by an effective enumeration of $S$ we mean an enumeration $S=\{x_{i}\}_{i\in\mathbb{N}}$ satisfying that there exists an algorithm $\mathcal{A}$ which, upon input $i\in\mathbb{N}$ outputs $x_{i}$.
Definition 3.2.
A real number $x$ is called
  1. (i)
    computable if there exist two computable functions $f\colon\mathbb{N}\rightarrow\mathbb{Z}$ and $g\colon\mathbb{N}\rightarrow\mathbb{N}$ satisfying that for each $n\in\mathbb{N}$, $\bigl{\lvert}\frac{f(n)}{g(n)}-x\bigr{\rvert}<2^{-n}.$
  2. (ii)
    lower-computable if there exist two computable functions $f\colon\mathbb{N}\rightarrow\mathbb{Z}$ and $g\colon\mathbb{N}\rightarrow\mathbb{N}$ satisfying that $\bigl{\{}\frac{f(n)}{g(n)}\bigr{\}}_{n\in\mathbb{N}}$ is an increasing sequence and converges to $x$ as $n\rightarrow+\infty$.
  3. (iii)
    upper-computable if there exist two computable functions $f\colon\mathbb{N}\rightarrow\mathbb{Z}$ and $g\colon\mathbb{N}\rightarrow\mathbb{N}$ satisfying that $\bigl{\{}\frac{f(n)}{g(n)}\bigr{\}}_{n\in\mathbb{N}}$ is a decreasing sequence and converges to $x$ as $n\rightarrow+\infty$.
3.2. Computable metric spaces
The above definitions equip the real numbers with a computability structure. This can be extended to virtually any separable metric space. We now give a short introduction.
Definition 3.3.
A computable metric space is a triple $(X,\,\rho,\,\mathcal{S})$, where
  1. (i)
    $(X,\rho)$ is a separable metric space;
  2. (ii)
    $\mathcal{S}=\{s_{n}:n\in\mathbb{N}\}$ is a dense subset of $X$;
  3. (iii)
    there exists an algorithm which, on input $i,\,j,\,m\in\mathbb{N}$, outputs $y_{i,j,m}\in\mathbb{Q}$ satisfying $\lvert y_{i,j,m}-\rho(s_{i},s_{j})\rvert<2^{-m}$.
The points in $\mathcal{S}$ are said to be ideal. Using the canonical bijection between $\mathbb{N}^{3}$ and $\mathbb{N}$ (see Subsection 2.1), we fix the effective enumeration of ideal balls $\{B_{i}\}_{i\in\mathbb{N}}$ for the set $\{B(s_{i},j/k):i,\,j,\,k\in\mathbb{N}\}$ of balls with rational radii centered at points in $\mathcal{S}$. These balls are called the ideal balls in $(X,\,\rho,\,\mathcal{S})$.
Definition 3.4.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. We say that a point $x\in X$ is computable if there exists a computable function $f\colon\mathbb{N}\rightarrow\mathbb{N}$ such that $\rho(s_{f(n)},x)<2^{-n}$ for each $n\in\mathbb{N}.$ Moreover, a sequence of points $\{x_{i}\}_{i\in\mathbb{N}}$ is said to be a uniformly computable sequence of points if there exists a computable function $f\colon\mathbb{N}^{2}\rightarrow\mathbb{N}$ such that $\rho(s_{f(n,m)},x_{m})<2^{-n}$ for all $n,\,m\in\mathbb{N}$.
As a remark, a finite sequence of computable points is always a uniformly computable sequence of points. Similarly, for other definitions of computable objects detailed below, we always have that a finite sequence of computable objects is always a uniformly computable sequence.
Definition 3.5.
In a computable metric space $(X,\,\rho,\,\mathcal{S})$, an open set $U\subseteq X$ is called lower-computable if there is a computable function $f\colon\mathbb{N}\rightarrow\mathbb{N}$ such that $U=\bigcup_{n\in\mathbb{N}}B_{f(n)}$. A closed set $K$ is said to be upper-computable if its complement is a lower-computable open set.
Definition 3.6.
In a computable metric space $(X,\,\rho,\,\mathcal{S})$, a family $\{U_{i}\}_{i\in\mathbb{N}}$ of lower-computable open sets is called a uniformly computable sequence of lower-computable open sets if there is a computable function $f\colon\mathbb{N}^{2}\rightarrow\mathbb{N}$ such that $U_{i}=\bigcup_{n\in\mathbb{N}}B_{f(i,n)}$.
As a consequence of Definitions 3.5 and 3.6, the canonical bijection between $\mathbb{N}^{2}$ and $\mathbb{N}$ (see Subsection 2.1) gives the following proposition.
Proposition 3.7.
For a uniformly computable sequence $\{U_{i}\}_{i\in\mathbb{N}}$ of lower-computable open sets in a computable metric space $(X,\,\rho,\,\mathcal{S})$, their union $U=\bigcup_{i\in\mathbb{N}}U_{i}$ is a lower-computable open set.
Before the definition of computable functions between computable metric spaces, we recall the definition of an oracle of a point in computable metric space.
Definition 3.8.
Given a computable metric space $(X,\,\rho,\,\mathcal{S})$, and a point $x\in X$, we say that a function $\varphi\colon\mathbb{N}\rightarrow\mathbb{N}$ is an oracle for $x\in X$ if $\rho(s_{\varphi(m)},x)<2^{-m}$ for each $m\in\mathbb{N}.$
Definition 3.9.
Assume that $(X,\,\rho,\,\mathcal{S})$ and $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$ are two computable metric spaces. Then a function $f\colon X\rightarrow X^{\prime}$ is computable if there exists an algorithm which, for each $x\in X$ and each $n\in\mathbb{N}$, on input $n\in\mathbb{N}$ and an oracle $\varphi$ for $x$, outputs $m\in\mathbb{N}$ satisfying that $\rho^{\prime}(s^{\prime}_{m},f(x))<2^{-n}$. Moreover, a sequence $\{f_{i}\}_{i\in\mathbb{N}}$ of functions $f_{i}\colon X\rightarrow X^{\prime}$ is called a uniformly computable sequence of functions if there exists an algorithm which, for each $x\in X$ and each $n\in\mathbb{N}$, on input $i,\,n\in\mathbb{N}$, and an oracle $\varphi$ for $x$, outputs $m\in\mathbb{N}$ satisfying that $\rho^{\prime}(s^{\prime}_{m},f_{i}(x))<2^{-n}$.
For example, in [Weih00], Examples 4.3.3 and 4.3.13.5 give the computability of the exponential function $\exp\colon\mathbb{R}\rightarrow\mathbb{R}$ and the logarithmic function $\log\colon\mathbb{R}^{+}\rightarrow\mathbb{R}$, respectively. The following proposition is a classical result that describes a topological property of a computable function.
Proposition 3.10.
Assume that $(X,\,\rho,\,\mathcal{S})$ and $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$ are two computable metric spaces with $\mathcal{S}=\{s_{i}\}_{i\in\mathbb{N}}$ and $\mathcal{S}^{\prime}=\{s_{i}^{\prime}\}_{i\in\mathbb{N}}$, and $\{B_{i}^{\prime}\}_{i\in\mathbb{N}}$ is the effective enumeration of ideal balls of $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$. If $f\colon X\rightarrow X^{\prime}$ is computable, then $\bigl{\{}f^{-1}(B_{i}^{\prime})\bigr{\}}_{i\in\mathbb{N}}$ is a uniformly computable sequence of lower-computable open sets in the computable metric space $(X,\,\rho,\,\mathcal{S})$, and preimages of lower-computable open sets are still lower-computable open sets.
Proof.
Assume that $f\colon X\rightarrow X^{\prime}$ is a computable function. First, we demonstrate that $\bigl{\{}f^{-1}(B_{i}^{\prime})\bigr{\}}_{i\in\mathbb{N}}$ is a uniformly computable sequence of lower-computable open sets in the computable metric space $(X,\,\rho,\,\mathcal{S})$. Fix an integer $i\in\mathbb{N}$. Now we compute $\{c_{j}\}_{j\in\mathbb{N}}$ and $\{r_{j}\}_{j\in\mathbb{N}}$ satisfying that $f^{-1}(B_{i}^{\prime})=\bigcup_{j\in\mathbb{N}}B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}$. Since $\{B_{i}^{\prime}\}_{i\in\mathbb{N}}$ is an effective enumeration, we can compute $m\in\mathbb{N}$ and $r\in\mathbb{Q}$ satisfying that $B_{\rho^{\prime}}(s_{m}^{\prime},r)=B_{i}^{\prime}$. Then we say that a sequence $\{p_{i}\}_{i=1}^{q}$ of integers is admissible if $B_{\rho}\bigl{(}s_{p_{i+1}},2^{-i-1}\bigr{)}\subseteq B_{\rho}\bigl{(}s_{p_{i}},2^{-i}\bigr{)}$ for each $i\in\mathbb{N}$. Note that $(X,\,\rho,\,\mathcal{S})$ is a computable metric space. It follows from statement (iii) of Definition 3.3 that we can compute all the possible admissible sequences. We enumerate them by $\{P_{i}\}_{i\in\mathbb{N}}$ effectively. Moreover, for each admissible sequence $P=\{p_{i}\}_{i=1}^{q}$, the function $\varphi_{P}\colon\mathbb{N}\rightarrow\mathbb{N}$ given by
$\varphi_{P}(i)\coloneqq\begin{cases}p_{i}&\text{if }1\leqslant i\leqslant q;\\ p_{q}&\text{if }i\geqslant q+1\end{cases}$
is an oracle for $s_{p_{q}}$. Since $f$ is computable, there exists an algorithm $M_{f}$ which, for each $x\in X$ and each $n\in\mathbb{N}$, on input $n\in\mathbb{N}$ and an oracle $\varphi$ for $x$, outputs $l\in\mathbb{N}$ satisfying that $\rho^{\prime}(s^{\prime}_{l},f(x))<2^{-n}$. Now we design the algorithm $M_{f}^{\prime}$ to compute $\{c_{j}\}_{j\in\mathbb{N}}$ and $\{r_{j}\}_{j\in\mathbb{N}}$ satisfying that $f^{-1}(B_{i}^{\prime})=\bigcup_{j\in\mathbb{N}}B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}$ as follows. We enumerate $\mathbb{N}\times\mathbb{N}$ as $\{(i_{u},n_{u})\}_{u\in\mathbb{N}}$ effectively.
  1. (i)
    Set $u$ to be $1$ and $j$ to be $1$.
  2. (ii)
    We input the oracle $\varphi_{P_{i_{u}}}$ and $n_{u}$ in $M_{f}$ to obtain the output $m_{i_{u},n_{u}}$. Assume that $M_{f}$ only reads $\{\varphi_{P_{i_{u}}}(v):v\in[1,k_{i_{u},n_{u}}]\cap\mathbb{N}\}$ to obtain this output.
  3. (iii)
    We check whether $B_{\rho^{\prime}}\bigl{(}s^{\prime}_{m_{i_{u},n_{u}}},2^{-n_{u}}\bigr{)}\subseteq B_{\rho^{\prime}}(s^{\prime}_{m},r)$. If so, then define $c_{j}\coloneqq\varphi_{P_{i_{u}}}(k_{i_{u},n_{u}})$ and $r_{j}\coloneqq 2^{-k_{i_{u},n_{u}}}$, and set $j$ to be $j+1$.
  4. (iv)
    Set $u$ to be $u+1$ and come back to Step (ii).
Now we verify that $f^{-1}(B_{i}^{\prime})=\bigcup_{j\in\mathbb{N}}B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}$. Fix an integer $j\in\mathbb{N}$. Now we demonstrate that $f\bigl{(}B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}\bigr{)}\subseteq B_{i}^{\prime}$. By Step (iii) of $M_{f}^{\prime}$, we have $c_{j}=\varphi_{P_{i_{u}}}(k_{i_{u},n_{u}})$ and $r_{j}=2^{-k_{i_{u},n_{u}}}$ for some $u\in\mathbb{N}$ with $B_{\rho^{\prime}}\bigl{(}s^{\prime}_{m_{i_{u},n_{u}}},2^{-n_{u}}\bigr{)}\subseteq B_{\rho^{\prime}}(s^{\prime}_{m},r)$. For each $x\in B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}$, there exists an oracle $\widetilde{\varphi}_{P_{i_{u}},x}$ for $x$ satisfying that $\widetilde{\varphi}_{P_{i_{u}},x}(v)=\varphi_{P_{i_{u}}}(v)$ for each integer $v\in[1,k_{i_{u},n_{u}}]$. If we input $\widetilde{\varphi}_{P_{i_{u}},x}$ and $n_{u}$ in $M_{f}$, then the output is still $m_{i_{u},n_{u}}$. Then it follows from the definition of $M_{f}$ and Step (ii) of $M_{f}^{\prime}$ that $f(x)\in B_{\rho^{\prime}}\bigl{(}s^{\prime}_{m_{i_{u},n_{u}}},2^{-n_{u}}\bigr{)}$. To sum up, we have $f\bigl{(}B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}\bigr{)}\subseteq B_{\rho^{\prime}}\bigl{(}s^{\prime}_{m_{i_{u},n_{u}}},2^{-n_{u}}\bigr{)}\subseteq B_{\rho^{\prime}}(s^{\prime}_{m},r)=B_{i}^{\prime}$. Thus, $f^{-1}(B_{i}^{\prime})\supseteq\bigcup_{j\in\mathbb{N}}B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}$. Next, we establish that $f^{-1}(B_{i}^{\prime})\subseteq\bigcup_{j\in\mathbb{N}}B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}$. Fix an arbitrary point $x\in f^{-1}(B_{i}^{\prime})$, it suffices to check $x\in B_{\rho}\bigl{(}s_{c_{j}},r_{j}\bigr{)}$ for some $j\in\mathbb{N}$. Since $\mathcal{S}$ is dense in $X$, there exists an oracle $\overline{\varphi}_{x}$ for $x$ satisfying that $\{\overline{\varphi}_{x}(v)\}_{v=1}^{q}$ is an admissible sequence for each $q\in\mathbb{N}$. Since $f(x)\in B_{i}^{\prime}$, there exists $n(x)\in\mathbb{N}$ satisfying that $B_{\rho^{\prime}}\bigl{(}f(x),2^{-n(x)+1}\bigr{)}\subseteq B_{i}^{\prime}$. We input $\overline{\varphi}_{x}$ and $n(x)$ in $M_{f}$ to obtain the output $m(x)$, then $\rho^{\prime}\bigl{(}s^{\prime}_{m(x)},f(x)\bigr{)}<2^{-n(x)}$, hence, $B_{\rho^{\prime}}\bigl{(}s^{\prime}_{m(x)},2^{-n(x)}\bigr{)}\subseteq B_{\rho^{\prime}}\bigl{(}f(x),2^{-n(x)+1}\bigr{)}\subseteq B_{i}^{\prime}$. Assume that $M_{f}$ only reads $\{\overline{\varphi}_{x}(v):v\in[1,k(x)]\cap\mathbb{N}\}$ to obtain this output. Note that $Q=\{\overline{\varphi}_{x}(v)\}_{v=1}^{k(x)}$ is an admissible sequence. Then in Step (ii) of $M_{f}^{\prime}$, we will input $\varphi_{Q}$ and $n(x)$ in $M_{f}$. It follows from the definition of $k(x)$ and $Q$ that this output is still $m(x)$. Since $B_{\rho^{\prime}}\bigl{(}s^{\prime}_{m(x)},2^{-n(x)}\bigr{)}\subseteq B_{i}^{\prime}$, it follows from the definition of $\varphi_{Q}$ that $c_{j(x)}=\varphi_{Q}(k(x))=\overline{\varphi}_{x}(k(x))$ and $r_{j(x)}=2^{-k(x)}$ for some $j(x)\in\mathbb{N}$. By the definition of $\overline{\varphi}_{x}$, we have $x\in B_{\rho}\bigl{(}s_{\varphi_{x}(k(x))},2^{-k(x)}\bigr{)}=B_{\rho}\bigl{(}s_{c_{j(x)}},r_{j(x)}\bigr{)}$. To sum up, one can see that $f^{-1}(B_{i}^{\prime})=\bigcup_{j\in\mathbb{N}}B(s_{c_{j}},r_{j})$. Hence, by Definition 3.6, $\bigl{\{}f^{-1}(B_{i}^{\prime})\bigr{\}}_{i\in\mathbb{N}}$ is a uniformly computable sequence of lower-computable open sets in the computable metric space $(X,\,\rho,\,\mathcal{S})$. Therefore, it follows from Definition 3.5 and Proposition 3.7 that preimages of lower-computable open sets are still lower-computable open sets. ∎
According to Definition 3.9, we can obtain the following corollary.
Corollary 3.11.
Assume that $(X,\,\rho,\,\mathcal{S})$ and $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$ are two computable metric spaces with $\mathcal{S}=\{s_{i}\}_{i\in\mathbb{N}}$ and $\mathcal{S}^{\prime}=\{s_{i}^{\prime}\}_{i\in\mathbb{N}}$, and $\{B_{i}^{\prime}\}_{i\in\mathbb{N}}$ is the effective enumeration of ideal balls of $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$. If a sequence $\{f_{i}\}_{i\in\mathbb{N}}$ of functions $f_{i}\colon X\rightarrow X^{\prime}$ is a uniformly computable sequence of functions, then $\bigl{\{}f_{j}^{-1}(B_{i}^{\prime}):i,\,j\in\mathbb{N}\bigr{\}}$ is a uniformly computable sequence of lower-computable open sets in the computable metric space $(X,\,\rho,\,\mathcal{S})$.
Definition 3.12.
Fix an effective enumeration $\{q_{n}\}_{n\in\mathbb{N}}$ of $\mathbb{Q}$. Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $C$ be a subset of $X$. A function $T\colon X\rightarrow\mathbb{R}$ is said to be upper-computable on $C$ if there is a uniformly computable sequence $\{U_{i}\}_{i\in\mathbb{N}}$ of lower-computable open sets in the computable metric space $(X,\,\rho,\,\mathcal{S})$ satisfying that
$T^{-1}(-\infty,q_{i})\cap C=U_{i}\cap C.$
The set $C$ is called the domain of computability of $T$.
Definition 3.13.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $T\colon X\rightarrow X$ a function with $p=p(x)\coloneqq\operatorname{card}\bigl{(}T^{-1}(x)\bigr{)}<+\infty$ for each $x\in X$. Then the inverse $T^{-1}$ of $T$ is said to be computable if there exists an algorithm $\mathcal{A}$ such that for each $x\in X$ and each $n\in\mathbb{N}$, we can input an oracle $\varphi$ for $x$ in $\mathcal{A}$ to produce $\{y_{i}:1\leqslant i\leqslant p(x)\}$ that satisfies the following property: If $T^{-1}(x)=\{x_{i}:1\leqslant i\leqslant p(x)\}$, then for each $1\leqslant i\leqslant p(x)$, there exists $1\leqslant j\leqslant p(x)$ such that $\rho(x_{i},y_{j})<2^{-n}.$
It follows immediately from Definition 3.9 that the computable real-valued functions are closed under finitely many operations from the following list: addition, multiplication, division, scalar multiplication, $\max,$ and $\min$ (see for example, [Weih00, Corollary 4.3.4]). At the end of this subsection, we recall the definitions of a recursively compact set and a recursively precompact metric space introduced in [GHR10, Section 2].
Definition 3.14.
In a computable metric space $(X,\,\rho,\,\mathcal{S})$, a set $K\subseteq X$ is said to be recursively compact if it is compact and there is an algorithm which, on input $\{i_{j}\}_{j=1}^{p}\subseteq\mathbb{N}$ and $\{q_{j}\}_{j=1}^{p}\subseteq\mathbb{Q}^{+}$, halts if and only if $K\subseteq\bigcup_{j=1}^{p}B(s_{i_{j}},q_{j})$.
[GHR10, Proposition 1] gives some basic properties of recursively compact sets.
Proposition 3.15.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Then the following statements hold:
  1. (i)
    A point $x\in X$ is computable if and only if the singleton $\{x\}$ is a recursively compact set.
  2. (ii)
    Assume that $K\subseteq X$ is a recursively compact set, and $U\subseteq X$ is a lower-computable open set. Then $K\setminus U$ is recursively compact.
Definition 3.16.
A computable metric space $(X,\,\rho,\,\mathcal{S})$ is recursively precompact if there is an algorithm which, on input $n\in\mathbb{N}$, outputs a subset $\{i_{1},\,i_{2},\,\ldots,\,i_{p}\}$ of $\mathbb{N}$ satisfying that $\{s_{i_{1}},\,s_{i_{2}},\,\ldots,\,s_{i_{p}}\}$ is a $2^{-n}$-net of $X$.
3.3. Computability of probability measures
Following [HR09, Proposition 4.1.3], if $(X,\,\rho,\,\mathcal{S})$ is a computable metric space, and $X$ is bounded, then $\bigl{(}\mathcal{P}(X),\,W_{\rho},\,\mathcal{R}_{\mathcal{S}}\bigr{)}$ is a computable metric space, where the Wasserstein–Kantorovich metric $W_{\rho}$ is recalled in (2.2), and $\mathcal{R}_{\mathcal{S}}\subseteq\mathcal{P}(X)$ is the set of the probability measures such that each measure is supported on finitely many points of $\mathcal{S}$ and assigns rational values to them.
Definition 3.17.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $X$ be a bounded set. Then a computable measure $\mu$ is a computable point of $\bigl{(}\mathcal{P}(X),\,W_{\rho},\,\mathcal{R}_{\mathcal{S}}\bigr{)}$.
By [GHR10, Lemma 2.12 and Proposition 4], due to the completeness of $\mathcal{P}(X)$ with respect to the metric $W_{\rho}$, one can conclude the following result.
Proposition 3.18.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $X$ be recursively compact. Then $\mathcal{P}(X)$ is a recursively compact set in $\bigl{(}\mathcal{P}(X),\,W_{\rho},\,\mathcal{R}_{\mathcal{S}}\bigr{)}$.
Proposition 3.19 ([HR09, Corollary 4.3.2]).
Assume that $(X,\,\rho,\,\mathcal{S})$ is a computable metric space, and $\{f_{i}\}_{i\in\mathbb{N}}$ is a uniformly computable sequence of real-valued functions on $X$. If there is an algorithm which, on input $i\in\mathbb{N}$, outputs $M_{i}$ such that $\|f_{i}\|_{\infty}\leqslant M_{i}$, then the sequence of integral operators $\mathcal{I}_{i}\colon\mathcal{P}(X)\rightarrow\mathbb{R}$, $i\in\mathbb{N}$, defined by $\mathcal{I}_{i}(\mu)\coloneqq\int\!f_{i}\,\mathrm{d}\mu$, is a uniformly computable sequence of functions.
Assume, in addition, that $X$ is recursively compact, then by [BRY12, Proposition 2.13], upper bounds of computable functions on $X$ can be computed. Hence, we have the following result.
Corollary 3.20.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $X$ be recursively compact. Assume that $\{f_{i}\}_{i\in\mathbb{N}}$ is a uniformly computable sequence of real-valued functions on $X$. Then the sequence of integral operators $\mathcal{I}_{i}\colon\mathcal{P}(X)\rightarrow\mathbb{R}$, $i\in\mathbb{N}$, defined by $\mathcal{I}_{i}(\mu)\coloneqq\int\!f_{i}\,\mathrm{d}\mu$, is a uniformly computable sequence of functions.
Finally, we introduce a family of computable functions.
Definition 3.21.
Let $(X,\rho)$ be a metric space. Consider arbitrary constants $r,\,\epsilon>0$, and a point $u\in X$. Then the function $g_{u,r,\epsilon}$ given by
(3.1) $g_{u,r,\epsilon}(x)\coloneqq\biggl{\lvert}1-\frac{\lvert\rho(x,u)-r\rvert^{+}}{\epsilon}\biggr{\rvert}^{+}\quad\text{for each }x\in X,$
[…]
The hat function $g_{u,r,\epsilon}(x)$ is an $\epsilon^{-1}$-Lipschitz function which equals to $1$ in the ball $B(u,r)$, to 00 outside $B(u,r+\epsilon)$, and lies strictly between 00 and $1$ in $B(u,r+\epsilon)\setminus\overline{B(u,r)}$.
Definition 3.22.
[…]
According to the canonical bijection between $N^{*}$ and $N$ (see Subsection 2.1), it follows from Definitions 3.21 and 3.22 that there exists an enumeration $\{\varphi_{j}\}_{j\in\mathbb{N}}$ for $\mathcal{E}(\mathcal{S})$ such that $\{\varphi_{j}\}_{j\in\mathbb{N}}$ is a uniformly computable sequence of functions. We call such an enumeration a computable enumeration. Assume that $X$ is a compact separable metric space and $\mathcal{S}$ a dense subset of $X$, then it follows from the lattice version of the Stone–Weierstrass theorem that $\mathcal{E}(\mathcal{S})$ is dense in $C(X)$. Thus, using the dominated convergence theorem, we conclude the following proposition.
Proposition 3.23.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, $X$ be recursively compact, and $\{\varphi_{j}\}_{j\in\mathbb{N}}$ be a computable enumeration of $\mathcal{E}(\mathcal{S})$. Then for each pair of Borel probability measures $\mu,\,\nu\in\mathcal{P}(X)$, $\mu=\nu$ if and only if $\int\!\varphi_{j}\,\mathrm{d}\mu=\int\!\varphi_{j}\,\mathrm{d}\nu$, for each $j\in\mathbb{N}.$
4. Computable Analysis on thermodynamic formalism
In this section, we establish Theorems 1.1 and 1.2. In Subsection 4.1, we estimate the rates of convergences of iterations of some normalized Ruelle–Perron–Frobenius operators. In Subsection 4.2, we prove that the Jacobian function $J_{\varphi}(x)$ (see (4.19)) is computable. In Subsection 4.3, we demonstrate that two subsets $U$ and $W$ of $\mathcal{P}(X)$ (from Lemmas 4.12 and 4.13) are both lower-computable open sets. In Subsection 4.4, we complete the proofs of Theorems 1.1 and 1.2.
4.1. Ruelle–Perron–Frobenius operators and cones
In this subsection, we develop the technique of cones, combining ideas from [PU10], to prove in Theorem 4.1 that the convergence of iterations of some normalized Ruelle–Perron–Frobenius operator is at an exponential rate with explicit expressions for the related constants. The technique of cones first used by Birkhoff [Bi57] in the study of Banach spaces geometry was reintroduced by Liverani [Liv95a, Liv95b] for the analysis of the decay of correlations. For a quick introduction to cones in the context of connected spaces, see [Vi97, Section 2.2]. The lack of connectedness of $X$ becomes a key obstacle in our study of Computable Analysis. Due to this obstacle and the demand on the computability (in the sense of Computable Analysis) of various constants involved, we need to further refine the technique of cones. In this subsection, we denote
(4.1) $\widehat{K}(\lambda_{1},b,b^{\prime})\coloneqq 2\log\biggl{(}\frac{1+\lambda_{1}}{1-\lambda_{1}}\cdot\frac{b^{2}}{b-b^{\prime}}\biggr{)}>0$
for $1\leqslant b^{\prime}<b$ and $\lambda_{1}\in(0,1)$.
Theorem 4.1.
[…]
$\bigl{\|}\mathcal{L}^{mk}_{\overline{\varphi}}(\mathbbm{1})(x)-u_{\varphi}(x)\bigr{\|}_{\infty}\leqslant CeZ(1-\exp(-Z))^{k-1}$
holds for each $k\in\mathbb{N}$ satisfying $0\leqslant Z(1-\exp(-Z))^{k-1}\leqslant 1$, where $u_{\varphi}$ is the function from Corollary 2.9 and $Z\coloneqq\widehat{K}\bigl{(}\lambda_{1},C^{\prime},\frac{(C^{-2}Q^{m}+2)C^{\prime}}{2C^{-2}Q^{m}+2}\bigr{)}>0$.
The proof of Theorem 4.1 will be given at the end of this subsection. We begin by introducing some notations in the cone technique. Let $E$ be a vector space over $\mathbb{R}$. A convex cone in $E$ is a subset $\mathcal{C}\subseteq E\setminus\{0\}$ satisfying the following properties:
  1. (i)
    $tv\in\mathcal{C}$ for all $v\in\mathcal{C}$ and $t>0$.
  2. (ii)
    $t_{1}v_{1}+t_{2}v_{2}\in\mathcal{C}$ for all $v_{1},\,v_{2}\in\mathcal{C}$ and $t_{1},\,t_{2}>0$.
  3. (iii)
    $\overline{\mathcal{C}}\cap(-\overline{\mathcal{C}})=\{0\}$.
Here $\overline{\mathcal{C}}$ is the set consisting of $w\in E$ satisfying that there exist $v\in\mathcal{C}$ and $\{t_{n}\}_{n\in\mathbb{N}}\subseteq\mathbb{R}^{+}$ such that $w+t_{n}v\in\mathcal{C}$ for each $n\in\mathbb{N}$ and that $t_{n}$ converges to 00 as $n$ tends to +$\infty$+ ∞. Denote $-\overline{\mathcal{C}}\coloneqq\{-v\in E:v\in\overline{\mathcal{C}}\}$. Define
(4.2) $\alpha(v_{1},v_{2})\coloneqq\sup\{t>0:v_{2}-tv_{1}\in\mathcal{C}\}\quad\text{and}\quad\beta(v_{1},v_{2})\coloneqq\inf\{s>0:sv_{1}-v_{2}\in\mathcal{C}\}$
for all $v_{1},\,v_{2}\in\mathcal{C}$, with the conventions $\sup\emptyset=0,\,\inf\emptyset=+\infty$, and
(4.3) $\theta(v_{1},v_{2})\coloneqq\log(\beta(v_{1},v_{2}))-\log(\alpha(v_{1},v_{2})),$
with $\theta(v_{1},v_{2})=+\infty$ if $\alpha(v_{1},v_{2})=0$ or $\beta(v_{1},v_{2})=+\infty$. By the definition of a convex cone, we have $\alpha(v_{1},v_{2})\leqslant\beta(v_{1},v_{2}),\,\alpha(v_{1},v_{2})<+\infty,$ and $\beta(v_{1},v_{2})>0$ for all $v_{1},\,v_{2}\in\mathcal{C}$. Thus $\theta(v_{1},v_{2})$ is well-defined and takes value in $[0,+\infty]$. We call $\theta(\cdot,\cdot)$ the projective metric (on $\mathcal{C}$) associated to the convex cone $\mathcal{C}$.
Proposition 4.2 ([Vi97, Proposition 2.2]).
Let $\mathcal{C}$ be a convex cone. Then the projective metric $\theta$ has the following properties:
  1. (i)
    $\theta(v_{1},v_{2})=\theta(v_{2},v_{1})$ for all $v_{1},\,v_{2}\in\mathcal{C}$,
  2. (ii)
    $\theta(v_{1},v_{2})+\theta(v_{2},v_{3})\geqslant\theta(v_{1},v_{3})$ for all $v_{1},\,v_{2},\,v_{3}\in\mathcal{C}$,
  3. (iii)
    for all $v_{1},\,v_{2}\in\mathcal{C}$, $\theta(v_{1},v_{2})=0$ if and only if there exists $t>0$ such that $v_{1}=tv_{2}$.
[…]
Proposition 4.3 ([Vi97, Proposition 2.3]).
[…]
$\theta_{2}(L(v),L(w))\leqslant\bigl{(}1-e^{-U}\bigr{)}\cdot\theta_{1}(v,w).$
To prove Theorem 4.1, we introduce the following family of convex cones.
Definition 4.4.
Let $(X,\rho)$ be a compact metric space. Given constants $\xi,\,a,\,v>0$, and $b>1$, denote by $\mathcal{C}(a,v,b)\subseteq C(X)$ the convex cone consisting of continuous functions $u\colon X\rightarrow\mathbb{R}$ satisfying the following properties:
  1. (i)
    $u(x)>0$ for each $x\in X$,
  2. (ii)
    $u(x)\leqslant\exp(a\rho(x,y)^{v})\cdot u(y)$ for all $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$,
  3. (iii)
    $u(x)\leqslant bu(y)$ for each pair of $x,\,y\in X$,
and let $\alpha_{a,v,b}$ and $\beta_{a,v,b}$ be the corresponding functions associated to the convex cone $\mathcal{C}(a,v,b)$ defined in (4.2), and $\theta_{a,v,b}$ be the projective metric associated to the convex cone $\mathcal{C}(a,v,b)$ defined in (4.3).
Let us compute the expression of $\alpha_{a,v,b}$ for a given pair of $\varphi_{1},\varphi_{2}\in\mathcal{C}(a,v,b)$. By (4.2), $\alpha_{a,v,b}(\varphi_{1},\varphi_{2})$ is the supremum of all positive real numbers $t$ satisfying the following three properties for all $s,\,m,\,n,\,x,\,y\in X$ with $0<\rho(x,y)\leqslant\xi$:
  1. (i)
    $\varphi_{2}(s)-t\varphi_{1}(s)>0$, or equivalently, $t<\frac{\varphi_{2}(s)}{\varphi_{1}(s)}$;
  2. (ii)
    $\frac{\varphi_{2}(x)-t\varphi_{1}(x)}{\varphi_{2}(y)-t\varphi_{1}(y)}\leqslant\exp(a\rho(x,y)^{v})$, or equivalently, $t\leqslant\frac{e^{a\rho(x,y)^{v}}\varphi_{2}(y)-\varphi_{2}(x)}{e^{a\rho(x,y)^{v}}\varphi_{1}(y)-\varphi_{1}(x)}$;
  3. (iii)
    $\frac{\varphi_{2}(n)-t\varphi_{1}(n)}{\varphi_{2}(m)-t\varphi_{1}(m)}\leqslant b$, or equivalently, $t\leqslant\frac{b\varphi_{2}(m)-\varphi_{2}(n)}{b\varphi_{1}(m)-\varphi_{1}(n)}$.
In other words, $\alpha_{a,v,b}(\varphi_{1},\varphi_{2})$ is given by
(4.4) $\alpha_{a,v,b}(\varphi_{1},\varphi_{2})=\min\biggl{\{}\inf\limits_{x,\,y}\biggl{\{}\frac{e^{a\rho(x,y)^{v}}\varphi_{2}(x)-\varphi_{2}(y)}{e^{a\rho(x,y)^{v}}\varphi_{1}(x)-\varphi_{1}(y)}\biggr{\}},\,\inf\limits_{m,\,n\in X}\biggl{\{}\frac{b\varphi_{2}(m)-\varphi_{2}(n)}{b\varphi_{1}(m)-\varphi_{1}(n)}\biggr{\}},\,\inf\limits_{s\in X}\biggl{\{}\frac{\varphi_{2}(s)}{\varphi_{1}(s)}\biggr{\}}\biggr{\}},$
where the infimum in the first term is taken over all pairs of $x,\,y\in X$ with $0<\rho(x,y)\leqslant\xi$. We demonstrate that the images of the convex cone under iterations of the Ruelle–Perron–Frobenius operator contract uniformly.
Proposition 4.5.
Fix arbitrary constants $\eta,\,\xi,\,a_{0},\,v_{0}>0$, and $\lambda>1$. Let $(X,\rho)$ be a compact metric space, $T\colon X\rightarrow X$ a continuous, topologically exact, open map which is distance-expanding with respect to the metric $\rho$ with constants $\eta$, $\lambda$, and $\xi$, and the potential $\varphi\colon X\rightarrow\mathbb{R}$ a Hölder continuous function with respect to the metric $\rho$ with exponent $v_{0}$ and constant $a_{0}$. Put $\lambda_{1}\coloneqq\frac{1+\lambda^{-v_{0}}}{2}<1$ and $a^{\prime}\coloneqq\frac{6a_{0}\lambda^{v_{0}}-2a_{0}}{(\lambda^{v_{0}}-1)^{2}}$. Define $\overline{\varphi}(x)\coloneqq\varphi(x)-P(T,\varphi)$ for each $x\in X$, and $Q\coloneqq\exp(\inf_{x\in X}\overline{\varphi}(x))>0$. Fix a constant $b>2\exp(a^{\prime})$. Suppose that $\epsilon\in(0,\xi]$ is a constant with $2\exp(a^{\prime}\epsilon^{v_{0}})\leqslant b$, and that $G^{\prime}=\{x^{\prime}_{1},\,x^{\prime}_{2},\,\ldots,\,x^{\prime}_{n^{\prime}}\}$ is an $(\epsilon/2)$-net of $X$. Fix an arbitrary integer $m\in\mathbb{N}$ satisfying that $\bigcap_{i=1}^{n^{\prime}}T^{m}(B(x^{\prime}_{i},\epsilon/2))=X.$ Put $b^{\prime}\coloneqq\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2}<b$, where $C>1$ is from Lemma 2.8. Denote by $L\colon C(X)\rightarrow C(X)$ the operator given by
(4.5) $L(u)(x)\coloneqq\frac{\mathcal{L}_{\overline{\varphi}}(uu_{\varphi})(x)}{u_{\varphi}(x)}=\sum_{\overline{x}\in T^{-1}(x)}\frac{u(\overline{x})u_{\varphi}(\overline{x})}{u_{\varphi}(x)}\cdot\exp(\overline{\varphi}(\overline{x}))$
for all $u\in C(X)$ and $x\in X$. Here $u_{\varphi}$ is the function from Corollary 2.9. Then $L(\mathbbm{1})=\mathbbm{1}$ and
(4.6) $L^{m}(\mathcal{C}(a^{\prime},v_{0},b))\subseteq\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})\subseteq\mathcal{C}(a^{\prime},v_{0},b).$
Proof.
By (2.15) and (4.5), we have $L(\mathbbm{1})=\mathbbm{1}$. It remains to prove (4.6). Since $\lambda_{1}<1$ and $b^{\prime}<b$, it follows from Definition 4.4 that $\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})\subseteq\mathcal{C}(a^{\prime},v_{0},b)$. Fix an arbitrary $u\in\mathcal{C}(a^{\prime},v_{0},b)$. It suffices to demonstrate that $L^{m}(u)$ satisfies properties (i), (ii), and (iii) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})$. Since $u\in\mathcal{C}(a^{\prime},v_{0},b)$, then $u(x)>0$ for each $x\in X$. By (4.5), $L(u)(x)>0$ for each $x\in X$. By induction, we obtain that $L^{m}(u)(x)>0$ for each $x\in X$, so $L^{m}(u)$ satisfies property (i) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})$. Next, we prove that $L(u)$ satisfies property (ii) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})$, namely,
$L(u)(x)\leqslant\exp(\lambda_{1}a^{\prime}\rho(x,y)^{v_{0}})\cdot L(u)(y)$
for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$. Consider a pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$. Since $T$ is an open map which is distance-expanding with constants $\eta$, $\lambda$, and $\xi$, then for each $\overline{x}\in T^{-1}(x)$, one can find a point $y(\overline{x})\coloneqq T_{\overline{x}}^{-1}(y)\in T^{-1}(y)$ satisfying that $\rho(\overline{x},y(\overline{x}))\leqslant\lambda^{-1}\rho(x,y)\leqslant\xi,$ and hence,
(4.7) $u(\overline{x})\leqslant\exp(a^{\prime}\rho(\overline{x},y(\overline{x}))^{v_{0}})\cdot u(y(\overline{x}))\leqslant\exp(a^{\prime}\lambda^{-v_{0}}\rho(x,y)^{v_{0}})\cdot u(y(\overline{x})).$
Note that $\overline{\varphi}(x)=\varphi(x)-P(T,\varphi)$ is Hölder continuous with exponent $v_{0}$ and constant $a_{0}$. So for each $\overline{x}\in T^{-1}(x)$,
(4.8) $\exp(\overline{\varphi}(\overline{x}))\leqslant\exp(a_{0}\rho(\overline{x},y(\overline{x}))^{v_{0}})\cdot\exp(\overline{\varphi}(y(\overline{x})))\leqslant\exp(a_{0}\lambda^{-v_{0}}\rho(x,y)^{v_{0}})\cdot\exp(\overline{\varphi}(y(\overline{x}))).$
Moreover, by (2.16), for each $\overline{x}\in T^{-1}(x)$,
(4.9) $\frac{u_{\varphi}(\overline{x})}{u_{\varphi}(x)}\leqslant\exp\biggl{(}\frac{2a_{0}\rho(x,y)^{v_{0}}}{\lambda^{v_{0}}-1}\biggr{)}\cdot\frac{u_{\varphi}(y(\overline{x}))}{u_{\varphi}(y)}.$
Thus, by (4.5), (4.7), (4.8), and (4.9), we conclude that
$L(u)(x)$ $=\sum_{\overline{x}\in T^{-1}(x)}\frac{u(\overline{x})u_{\varphi}(\overline{x})}{u_{\varphi}(x)}\cdot\exp(\overline{\varphi}(\overline{x}))$
$\leqslant\sum_{\overline{x}\in T^{-1}(x)}\frac{u(y(\overline{x}))u_{\varphi}(y(\overline{x}))}{u_{\varphi}(y)}\cdot\exp\biggl{(}\overline{\varphi}(y(\overline{x}))+\biggl{(}\frac{a^{\prime}}{\lambda^{v_{0}}}+\frac{a_{0}}{\lambda^{v_{0}}}+\frac{2a_{0}}{\lambda^{v_{0}}-1}\biggr{)}\rho(x,y)^{v_{0}}\biggr{)}$
$=\exp(\lambda_{1}a^{\prime}\rho(x,y)^{v_{0}})\cdot L(u)(y)$
for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$, where the last equality holds due to the definitions of $\lambda_{1}$ and $a^{\prime}$. By induction, it follows from $\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})\subseteq\mathcal{C}(a^{\prime},v_{0},b)$ that $L^{m}(u)$ satisfies property (ii) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})$. Finally, we establish that $L^{m}(u)$ satisfies property (iii) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})$, namely,
(4.10) $L^{m}(u)(x)\leqslant b^{\prime}L^{m}(u)(y)\quad\text{for each pair of }x,\,y\in X.$
Without loss of generality, assume that $\inf_{x\in X}u(x)=1$. Let $x_{*}$ and $x^{*}$ be two points in $X$ satisfying that $u(x_{*})=1$ and $u(x^{*})=\sup_{x\in X}u(x)$. Since $u\in\mathcal{C}(a^{\prime},v_{0},b)$, we have $u(x^{*})\leqslant b$. First consider the case where $\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2}\leqslant u(x^{*})\leqslant b$. Consider a point $w\in X$. Since $G^{\prime}$ is an $(\epsilon/2)$-net of $X$, there exists an integer $i\in[1,n^{\prime}]$ with $x_{*}\in B(x^{\prime}_{i},\epsilon/2)$. Since $\bigcap_{i=1}^{n^{\prime}}T^{m}(B(x^{\prime}_{i},\epsilon/2))=X$, there exists $\overline{w}_{*}\in T^{-m}(w)$ such that $\rho(\overline{w}_{*},x^{\prime}_{i})<\epsilon/2$. Hence, $\rho(\overline{w}_{*},x_{*})\leqslant\rho(\overline{w}_{*},x^{\prime}_{i})+\rho(x^{\prime}_{i},x_{*})<\epsilon\leqslant\xi$. Moreover, since $u\in\mathcal{C}(a^{\prime},v_{0},b)$, we have
(4.11) $u(\overline{w}_{*})\leqslant\exp(a^{\prime}\rho(\overline{w}_{*},x_{*})^{v_{0}})\cdot u(x_{*})\leqslant\exp(a^{\prime}\epsilon^{v_{0}}).$
Hence, by the definitions of $\epsilon,\,Q$, $L(\mathbbm{1})=\mathbbm{1}$, (2.17), (4.5), and (4.11), we have
$L^{m}(u)(w)$ $=\frac{u(\overline{w}_{*})u_{\varphi}(\overline{w}_{*})}{u_{\varphi}(w)}\cdot e^{S_{m}\overline{\varphi}(\overline{w}_{*})}+\sum_{\overline{w}\in T^{-m}(w)\setminus\{\overline{w}_{*}\}}\frac{u(\overline{w})u_{\varphi}(\overline{w})}{u_{\varphi}(w)}\cdot e^{S_{m}\overline{\varphi}(\overline{w})}$
$\leqslant\frac{\exp(a^{\prime}\epsilon^{v_{0}})\cdot u_{\varphi}(\overline{w}_{*})}{u_{\varphi}(w)}\cdot e^{S_{m}\overline{\varphi}(\overline{w}_{*})}+u(x^{*})\sum_{\overline{w}\in T^{-m}(w)\setminus\{\overline{w}_{*}\}}\frac{u_{\varphi}(\overline{w})}{u_{\varphi}(w)}\cdot e^{S_{m}\overline{\varphi}(\overline{w})}$
$\leqslant u(x^{*})-(u(x^{*})-\exp(a^{\prime}\epsilon^{v_{0}}))\cdot C^{-2}\inf_{x\in X}\exp(S_{m}\overline{\varphi}(x))$
$\leqslant b-\biggl{(}\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2}-\frac{b}{2}\biggr{)}C^{-2}Q^{m}$
$=\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2},$
where the last equality comes from direct calculation. Next, we consider the case where $u(x^{*})\leqslant\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2}$. Since $L(\mathbbm{1})=\mathbbm{1}$, we have
$\sup_{w\in X}L^{m}(u)(w)\leqslant\sup_{w\in X}u(w)\leqslant\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2}.$
To sum up, we always have $\sup_{w\in X}L^{m}(u)(w)\leqslant\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2}$. Moreover, since $L(\mathbbm{1})=\mathbbm{1}$, it follows that $\inf_{w\in X}L^{m}(u)(w)\geqslant 1$. Therefore, we have estalished (4.10) for each pair of $x,\,y\in X$. ∎
Consider now the convex cone $\mathcal{C}_{+}$ of positive functions on $X$. Let $\alpha_{+}$ and $\beta_{+}$ be the corresponding functions associated to the convex cone $\mathcal{C}_{+}$ defined in (4.2). We denote by $\theta_{+}$ the projective metric associated to the convex cone $\mathcal{C}_{+}$ defined in (4.3). Now we establish the following result. Recall $\widehat{K}$ from (4.1).
Proposition 4.6.
For all constants $\xi,\,a,\,v>0$, $1<b^{\prime}<b$, and $\lambda_{1}\in(0,1)$, we have $\theta_{+}(v_{1},v_{2})=\log(\sup\{v_{2}(x)/v_{1}(x):x\in X\})-\log(\inf\{v_{2}(y)/v_{1}(y):y\in X\})\leqslant\theta_{a,v,b}(v_{1},v_{2})$ for each pair of $v_{1},\,v_{2}\in\mathcal{C}(a,v,b)$, and
(4.12) $\operatorname{diam}_{\theta_{a,v,b}}(\mathcal{C}(\lambda_{1}a,v,b^{\prime}))\leqslant\widehat{K}(\lambda_{1},b,b^{\prime}).$
Proof.
Consider a pair of $\varphi_{1},\,\varphi_{2}\in\mathcal{C}(\lambda_{1}a,v,b^{\prime})\subseteq\mathcal{C}_{+}$. By the definition of $\mathcal{C}(\lambda_{1}a,v,b^{\prime})$, we have
$\varphi_{2}(y)\leqslant\varphi_{2}(x)\exp(a\lambda_{1}\rho(x,y)^{v})\quad\text{and}\quad\varphi_{1}(y)\geqslant\varphi_{1}(x)\exp(-a\lambda_{1}\rho(x,y)^{v})$
for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$. Hence,
$\frac{\exp(a\rho(x,y)^{v})\varphi_{2}(x)-\varphi_{2}(y)}{\exp(a\rho(x,y)^{v})\varphi_{1}(x)-\varphi_{1}(y)}$ $\geqslant\frac{\varphi_{2}(x)}{\varphi_{1}(x)}\cdot\frac{\exp(a\rho(x,y)^{v})-\exp(a\lambda_{1}\rho(x,y)^{v})}{\exp(a\rho(x,y)^{v})-\exp(-a\lambda_{1}\rho(x,y)^{v})}$
(4.13) $\geqslant\inf\limits_{z>1}\left\{\frac{z-z^{\lambda_{1}}}{z-z^{-\lambda_{1}}}\right\}\cdot\frac{\varphi_{2}(x)}{\varphi_{1}(x)}$
$\geqslant\frac{1-\lambda_{1}}{1+\lambda_{1}}\cdot\frac{\varphi_{2}(x)}{\varphi_{1}(x)}$
for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$. By Definition 4.4, we have
$\varphi_{2}(y)\leqslant b^{\prime}\varphi_{2}(x)<b\varphi_{2}(x)\quad\text{and}\quad 0<\varphi_{1}(y)<b\varphi_{1}(x)$
for each pair of $x,\,y\in X$. Hence,
(4.14) $\frac{b\varphi_{2}(x)-\varphi_{2}(y)}{b\varphi_{1}(x)-\varphi_{1}(y)}\geqslant\frac{b-b^{\prime}}{b}\cdot\frac{\varphi_{2}(x)}{\varphi_{1}(x)}$
for each pair of $x,\,y\in X$. Consider an arbitrary pair of $v_{1},\,v_{2}\in\mathcal{C}(a,v,b)$. By the definition of $\alpha_{+}$ and $\beta_{+}$, we have $\alpha_{+}(v_{1},v_{2})=\inf_{t\in X}(v_{2}(t)/v_{1}(t))$, and $\beta_{+}(v_{1},v_{2})=\sup_{s\in X}(v_{2}(s)/v_{1}(s))$. Hence, $\theta_{+}(v_{1},v_{2})=\log(\sup\{v_{2}(x)/v_{1}(x):x\in X\})-\log(\inf\{v_{2}(y)/v_{1}(y):y\in X\}).$ By (4.4), we have $\alpha_{+}(v_{1},v_{2})\geqslant\alpha_{a,v,b}(v_{1},v_{2})$. Similarly, we have $\beta_{+}(v_{1},v_{2})\leqslant\beta_{a,v,b}(v_{1},v_{2})$. Therefore, $\theta_{+}(v_{1},v_{2})\leqslant\theta_{a,v,b}(v_{1},v_{2})$. Consider an arbitrary pair of $\varphi_{1},\,\varphi_{2}\in\mathcal{C}(\lambda_{1}a,v,b^{\prime})$. It follows from (4.4), (4.13), and (4.14) that $\alpha_{a,v,b}(\varphi_{1},\varphi_{2})\geqslant\frac{1-\lambda_{1}}{1+\lambda_{1}}\cdot\frac{b-b^{\prime}}{b}\cdot\alpha_{+}(\varphi_{1},\varphi_{2})$. Similarly, we can obtain that $\beta_{a,v,b}(\varphi_{1},\varphi_{2})\leqslant\frac{1+\lambda_{1}}{1-\lambda_{1}}\cdot\frac{b}{b-b^{\prime}}\cdot\beta_{+}(\varphi_{1},\varphi_{2})$. Moreover, by the definitions of $\theta_{a,v,b}$ and $\theta_{+}$, it follows from (4.1) that
$\theta_{a,v,b}(\varphi_{1},\varphi_{2})$ $=\log(\beta_{a,v,b}(\varphi_{1},\varphi_{2}))-\log(\alpha_{a,v,b}(\varphi_{1},\varphi_{2}))$
$\leqslant\log(\beta_{+}(\varphi_{1},\varphi_{2}))-\log(\alpha_{+}(\varphi_{1},\varphi_{2}))+2\log\biggl{(}\frac{1+\lambda_{1}}{1-\lambda_{1}}\cdot\frac{b}{b-b^{\prime}}\biggr{)}$
$=\theta_{+}(\varphi_{1},\varphi_{2})+\widehat{K}(\lambda_{1},b,b^{\prime})$
for each pair of $\varphi_{1},\,\varphi_{2}\in\mathcal{C}(\lambda_{1}a,v,b^{\prime})$. This completes the proof of (4.12). ∎
We are now ready to establish Theorem 4.1.
Proof of Theorem 4.1.
[…]
$L^{m}(\mathcal{C}(a^{\prime},v_{0},C^{\prime}))\subseteq\mathcal{C}\biggl{(}\lambda_{1}a^{\prime},v_{0},\frac{(C^{-2}Q^{m}+2)C^{\prime}}{2C^{-2}Q^{m}+2}\biggr{)},$
where $L$ is defined in (4.5). Thus, it follows from Proposition 4.6 that
(4.15) $\operatorname{diam}_{\theta_{a^{\prime},v_{0},C^{\prime}}}(L^{m}(\mathcal{C}(a^{\prime},v_{0},C^{\prime})))$ $\leqslant\operatorname{diam}_{\theta_{a^{\prime},v_{0},C^{\prime}}}\biggl{(}\mathcal{C}\biggl{(}\lambda_{1}a^{\prime},v_{0},\frac{(C^{-2}Q^{m}+2)C^{\prime}}{2C^{-2}Q^{m}+2}\biggr{)}\biggr{)}$
$\leqslant\widehat{K}\biggl{(}\lambda_{1},C^{\prime},\frac{(C^{-2}Q^{m}+2)C^{\prime}}{2C^{-2}Q^{m}+2}\biggr{)}.$
Obviously, we have $\mathbbm{1}\in\mathcal{C}(a^{\prime},v_{0},C^{\prime})$. By (2.16) and (2.17), it follows from $a^{\prime}>\frac{a_{0}}{\lambda^{v_{0}}-1}$ and $C^{2}<C^{\prime}$ that $1/u_{\varphi}\in\mathcal{C}(a^{\prime},v_{0},C^{\prime})$. Then by Proposition 4.3 for $L^{m}\colon\mathcal{C}(a^{\prime},v_{0},C^{\prime})\rightarrow\mathcal{C}(a^{\prime},v_{0},C^{\prime})$ and (4.15), we have
(4.16) $\theta_{a^{\prime},v_{0},C^{\prime}}\bigl{(}L^{mk}(1/u_{\varphi}),\mathbbm{1}\bigr{)}$ $\leqslant(1-\exp(-Z))^{k-1}\cdot\theta_{a^{\prime},v_{0},C^{\prime}}\bigl{(}L^{m}(1/u_{\varphi}),\mathbbm{1}\bigr{)}$
$\leqslant(1-\exp(-Z))^{k-1}Z$
for each $k\in\mathbb{N}$, where $Z\coloneqq\widehat{K}\bigl{(}\lambda_{1},C^{\prime},\frac{(C^{-2}Q^{m}+2)C^{\prime}}{2C^{-2}Q^{m}+2}\bigr{)}$. By Proposition 4.6, we have
(4.17) $\theta_{+}\bigl{(}L^{mk}(1/u_{\varphi}),\mathbbm{1}\bigr{)}\leqslant\theta_{a^{\prime},v_{0},C^{\prime}}\bigl{(}L^{mk}(1/u_{\varphi}),\mathbbm{1}\bigr{)}.$
Then by (4.5), (4.16), and (4.17), one can see that
(4.18) $\theta_{+}\bigl{(}\mathcal{L}^{mk}_{\overline{\varphi}}(\mathbbm{1}),u_{\varphi}\bigr{)}=\theta_{+}\bigl{(}\mathcal{L}^{mk}_{\overline{\varphi}}(\mathbbm{1})/u_{\varphi},\mathbbm{1}\bigr{)}=\theta_{+}\bigl{(}L^{mk}(1/u_{\varphi}),\mathbbm{1}\bigr{)}\leqslant Z(1-\exp(-Z))^{k-1}.$
Since $L(\mathbbm{1})=\mathbbm{1}$, one can see that
$\sup_{x\in X}\frac{\mathcal{L}^{mk}_{\overline{\varphi}}(\mathbbm{1})(x)}{u_{\varphi}(x)}=\sup_{x\in X}\{L^{mk}(1/{u_{\varphi}(x)})\}\geqslant 1\geqslant\inf_{x\in X}\{L^{mk}(1/{u_{\varphi}(x)})\}=\inf_{x\in X}\frac{\mathcal{L}^{mk}_{\overline{\varphi}}(\mathbbm{1})(x)}{u_{\varphi}(x)}.$
Hence, by Proposition 4.6, (2.17) and (4.18), we have
$\bigl{\|}\mathcal{L}^{mk}_{\overline{\varphi}}(\mathbbm{1})(x)-u_{\varphi}(x)\bigr{\|}_{\infty}$ $\leqslant\bigl{\|}u_{\varphi}(x)\bigr{\|}_{\infty}\cdot\biggl{\|}\frac{\mathcal{L}^{mk}_{\overline{\varphi}}(\mathbbm{1})(x)}{u_{\varphi}(x)}-1\biggr{\|}_{\infty}$
$\leqslant C\bigl{(}\exp\bigl{(}Z(1-\exp(-Z))^{k-1}\bigr{)}-1\bigr{)}.$
Since for each $x\in[0,1]$, $e^{x}\leqslant ex+1,$ we have
$\bigl{\|}\mathcal{L}^{mk}_{\overline{\varphi}}(\mathbbm{1})(x)-u_{\varphi}(x)\bigr{\|}_{\infty}\leqslant CeZ(1-\exp(-Z))^{k-1}$
for each $k\in\mathbb{N}$ satisfying $0\leqslant Z(1-\exp(-Z))^{k-1}\leqslant 1$. The proof is, therefore, complete. ∎
4.2. Proof of the computability of the Jacobian
In this subsection, we aim to establish uniform computability of the Jacobian, as stated in the following theorem.
Theorem 4.7.
There exists an algorithm that satisfies the following property: For each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, on the following input
  1. (i)
    two algorithms computing the potential $\varphi$ and the map $T$, respectively,
  2. (ii)
    an algorithm outputting a net of $X$ with any given precision (in the sense of Definition 3.16),
  3. (iii)
    an oracle of a point $x\in X$,
  4. (iv)
    two rational constants $a_{0}$ and $v_{0}$ satisfying that $\varphi$ is Hölder continuous with exponent $v_{0}$ and constant $a_{0}$,
  5. (v)
    three rational constants $\eta$, $\lambda,$ and $\xi$ satisfying that $T$ is distance-expanding with these constants (in the sense of Definition 2.1),
  6. (vi)
    a constant $n\in\mathbb{N}$,
the algorithm outputs a rational $2^{-n}$-approximation for the value of $J_{\varphi}(x)$, where the function $J_{\varphi}$ is the function defined by
(4.19) $J_{\varphi}(x)\coloneqq\frac{u_{\varphi}(T(x))}{u_{\varphi}(x)}\exp(P(T,\varphi)-\varphi(x)),\qquad x\in X.$
The proof will be given at the end of this subsection. We begin with designing an algorithm that computes the function $\mathcal{L}_{\varphi}(\mathbbm{1})$.
Proposition 4.8.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, $X$ a recursively compact set, and $T\colon X\rightarrow X$ a computable distance-expanding map with respect to the metric $\rho$ with constants $\eta$, $\lambda$, and $\xi$. Then the inverse $T^{-1}$ of $T$ is computable (in the sense of Definition 3.13).
Proof.
Consider $x\in X$ and an oracle $\varphi$ for $x$. By Definition 3.8, $\{x\}=\bigcap_{n\in\mathbb{N}}\overline{B(s_{\varphi(n)},2^{-n})}$. Hence $T^{-1}(x)=\bigcap_{n\in\mathbb{N}}T^{-1}\bigl{(}\overline{B(s_{\varphi(n)},2^{-n})}\bigr{)}$. By Definition 3.6, we get that $\bigl{\{}\bigl{(}\overline{B(s_{\varphi(n)},2^{-n})}\bigr{)}^{c}\bigr{\}}_{n\in\mathbb{N}}$ is a uniformly computable sequence of lower-computable open sets. Since $T$ is computable, it follows from Proposition 3.10 that the sequence $\bigl{\{}T^{-1}\bigl{(}\bigl{(}\overline{B(s_{\varphi(n)},2^{-n})}\bigr{)}^{c}\bigr{)}\bigr{\}}_{n\in\mathbb{N}}$ is a uniformly computable sequence of lower-computable open sets. By Proposition 3.7, $\bigl{(}T^{-1}(x)\bigr{)}^{c}=\bigcup_{n\in\mathbb{N}}T^{-1}\bigl{(}\bigl{(}\overline{B(s_{\varphi(n)},2^{-n})}\bigr{)}^{c}\bigr{)}$ is a lower-computable open set. By Proposition 3.15, noting that $X$ is recursively compact, $T^{-1}(x)$ is a recursively compact set. By [GHR10, Proposition 4], $X$ is recursively precompact, and consequently, there is an algorithm that outputs an $\eta$-net of $X$. Because $T$ is distance-expanding with the constants $\eta$, $\lambda$, and $\xi$, there is at most one preimage point of $x$ in each ball with radius $\eta$. Therefore each preimage point of $x$ is computable. ∎
As an immediate consequence of Proposition 4.8 and the computability of the exponential function, one gets the computability of the Ruelle–Perron–Frobenius operator in the following sense:
Corollary 4.9.
There exists an algorithm that satisfies the following property: For each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, this algorithm outputs a rational $2^{-n}$-approximation for the value of $\mathcal{L}^{m}_{\varphi}(\mathbbm{1})(x)$, after inputting the following data in this algorithm:
  1. (i)
    two algorithms computing the potential $\varphi$ and the transformation $T$, respectively,
  2. (ii)
    an algorithm outputting a net of $X$ with any given precision (in the sense of Definition 3.16),
  3. (iii)
    an oracle of a point $x\in X$,
  4. (iv)
    three rational constants $\eta$, $\lambda,$ and $\xi$ satisfying that $T$ is distance-expanding with these constants (in the sense of Definition 2.1),
  5. (v)
    two constants $n,\,m\in\mathbb{N}$.
We can now apply Theorem 4.1 to prove the computability of the pressure $P(T,\varphi)$ and the function $u_{\varphi}(x)$ (see Corollary 2.9 for the definition of the function $u_{\varphi}(x)$).
Lemma 4.10.
There exists an algorithm that satisfies the following property: For each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, this algorithm outputs a rational $2^{-n}$-approximation for the topological pressure $P(T,\varphi)$, after inputting the following data in this algorithm:
  1. (i)
    two algorithms computing the potential $\varphi$ and the map $T$, respectively,
  2. (ii)
    an algorithm outputting a net of $X$ with any given precision (in the sense of Definition 3.16),
  3. (iii)
    two rational constants $a_{0}$ and $v_{0}$ satisfying that $\varphi$ is Hölder continuous with exponent $v_{0}$ and constant $a_{0}$,
  4. (iv)
    three rational constants $\eta$, $\lambda,$ and $\xi$ satisfying that $T$ is distance-expanding with these constants (in the sense of Definition 2.1),
  5. (v)
    a constant $n\in\mathbb{N}$.
Proof.
We can design the algorithm following the steps below:
  1. (1)
    Compute a $\xi^{\prime}$-net $G$ for the space $X$, where $\xi^{\prime}\coloneqq\min\{\eta,\,\xi\}$. Then compute a constant $N\in\mathbb{N}$ satisfying that $\bigcap_{x\in G}\bigcup_{k=0}^{N}T^{k}(x,\xi)=X.$
  2. (2)
    Use [BRY12, Proposition 2.13] to compute $\left\|\varphi\right\|_{\infty}$.
  3. (3)
    Find $N_{1}\in\mathbb{N}$ with $2^{n+1}N\log(\operatorname{card}G)\cdot\bigl{(}\frac{4a_{0}\xi^{v_{0}}}{\lambda^{v_{0}}-1}+2N\|\varphi\|_{\infty}\bigr{)}\leqslant N_{1}$.
  4. (4)
    By Corollary 4.9, we compute and output the value of
    $v_{n}\approx w_{n}\coloneqq N_{1}^{-1}\log\bigl{(}\mathcal{L}_{\varphi}^{N_{1}}(\mathbbm{1})(x_{0})\bigr{)}$
    with precision $2^{-n-1}$, where $x_{0}\in\mathcal{S}$ is an ideal point.
Let us verify that $v_{n}$ satisfies $\lvert v_{n}-P(T,\varphi)\rvert<2^{-n}$ for each $n\in\mathbb{N}$. To see this, it suffices to check that $\lvert w_{n}-P(T,\varphi)\rvert<2^{-n-1}$ for each $n\in\mathbb{N}$. Since $T\colon X\rightarrow X$ is distance-expanding with constants $\eta$, $\lambda,$ and $\xi$, then $T$ is injective on each balls of radius $\xi^{\prime}_{0}$. Hence, we have $D\coloneqq\max_{x\in X}\operatorname{card}\bigl{(}T^{-1}(x)\bigr{)}\leqslant\operatorname{card}G$. Moreover, by Lemma 2.8, and steps (4) and (5) above, one can conclude that
$\lvert w_{n}-P(T,\varphi)\rvert=\bigl{\lvert}N_{1}^{-1}\log\bigl{(}e^{-{N_{1}}P(T,\varphi)}\mathcal{L}_{\varphi}^{N_{1}}(\mathbbm{1})(x_{0})\bigr{)}\bigr{\rvert}<(\log C)/N_{1}<2^{-n-1}.\qed$
Lemma 4.11.
There exists an algorithm that satisfies the following property: For each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, this algorithm outputs a rational $2^{-n}$-approximation for the value of $u_{\varphi}(x)$, after inputting the following data in this algorithm:
  1. (i)
    two algorithms computing the potential $\varphi$ and the map $T$, respectively,
  2. (ii)
    an algorithm outputting a net of $X$ with any given precision (in the sense of Definition 3.16),
  3. (iii)
    two rational constants $a_{0}$ and $v_{0}$ satisfying that $\varphi$ is Hölder continuous with exponent $v_{0}$ and constant $a_{0}$,
  4. (iv)
    three rational constants $\eta$, $\lambda,$ and $\xi$ satisfying that $T$ is distance-expanding with these constants (in the sense of Definition 2.1),
  5. (v)
    an oracle of a point $x\in X$,
  6. (vi)
    a constant $n\in\mathbb{N}$.
Proof.
We can design the algorithm following the steps below:
  1. (1)
    Compute a $\xi^{\prime}$-net $G$ for the space $X$, where $\xi^{\prime}\coloneqq\min\{\eta,\,\xi\}$. Then compute a constant $N\in\mathbb{N}$ satisfying that $\bigcap_{x\in G}\bigcup_{k=0}^{N}T^{k}(x,\xi)=X.$
  2. (2)
    Use Lemma 4.10 to compute the topological pressure $P(T,\varphi)$.
  3. (3)
    Use [BRY12, Proposition 2.13] to compute $\left\|\varphi\right\|_{\infty}$ and $Q\coloneqq\inf_{x\in X}\{\varphi(x)\}-P(T,\varphi)$.
  4. (4)
[…]
    $\overline{C}\coloneqq(\operatorname{card}G)^{N}\exp\Bigl{(}\frac{4a_{0}\xi^{v_{0}}}{\lambda^{v_{0}}-1}+2N\|\varphi\|_{\infty}\Bigr{)}.$
  5. (5)
    Use Theorem 4.1 to compute $\epsilon\in(0,\xi)$ with $2\exp(a^{\prime}\epsilon^{v_{0}})<C^{\prime}.$ Then compute a $(\epsilon/2)$-net $G^{\prime}$ for the space $X$, and a constant $m\in\mathbb{N}$ satisfying that $\bigcap_{t\in G^{\prime}}T^{m}(t,\epsilon/2)=X$.
  6. (6)
    Use (4.1) to compute $\overline{Z}\coloneqq\widehat{K}\Bigl{(}\frac{\lambda^{-v_{0}}+1}{2},C^{\prime},\frac{(\overline{C}^{-2}Q^{m}+2)C^{\prime}}{2\overline{C}^{-2}Q^{m}+2}\Bigr{)}$.
  7. (7)
    Find $k\in\mathbb{N}$ satisfying that
    (4.20) $0<\overline{Z}(1-\exp(-\overline{Z}))^{k-1}\leqslant 1\quad\text{and}\quad\overline{C}e\overline{Z}(1-\exp(-\overline{Z}))^{k-1}<2^{-n-1}.$
  8. (8)
    By Proposition 4.8 and Lemma 4.10, we compute and output the value of
    $s_{n}\approx t_{n}\coloneqq e^{-mkP(T,\varphi)}\mathcal{L}_{\varphi}^{mk}(\mathbbm{1})(x)$
    with precision $2^{-n-1}$.
[…]
$\lvert t_{n}-u_{\varphi}(x)\rvert\leqslant\bigl{\|}\mathcal{L}_{\overline{\varphi}}^{mk}(\mathbbm{1})(x)-u_{\varphi}(x)\bigr{\|}_{\infty}\leqslant CeZ\bigl{(}1-e^{-Z}\bigr{)}^{k-1}<\overline{C}e\overline{Z}\bigl{(}1-e^{-\overline{Z}}\bigr{)}^{k-1}<2^{-n-1},$
establishing the lemma. ∎
We are now ready to establish Theorem 4.7.
Proof of Theorem 4.7.
By Corollary 4.9, Lemmas 4.10, and 4.11, it follows from (4.19) that the function $J_{\varphi}$ is computable. ∎
4.3. Two subsets of $\mathcal{P}(X)$
In this subsection, we consider two subsets of $\mathcal{P}(X)$ closely related to the equilibrium state $\mu_{\varphi}$ and prove that they are both lower-computable open sets.
Lemma 4.12.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Assume that $X$ is recursively compact, and $T\colon X\rightarrow X$ is computable. Then the set $U$ of Borel probability measures that are not $T$-invariant is a lower-computable open set.
Proof.
Let $\{\varphi_{j}\}_{j\in\mathbb{N}}$ be a computable enumeration of $\mathcal{E}(\mathcal{S})$ (see Definition 3.22) in the computable metric space $(X,\,\rho,\,\mathcal{S})$. Claim. $U=\bigcup_{j\in\mathbb{N}}\Phi_{j}$, where $\Phi_{j}\coloneqq\left\{\mu\in\mathcal{P}(X):\int\!\varphi_{j}\,\mathrm{d}\mu-\int\!\varphi_{j}\circ T\,\mathrm{d}\mu\neq 0\right\}$ for each $j\in\mathbb{N}$. Indeed, consider $\mu\in\bigcap_{j\in\mathbb{N}}\Phi_{j}^{c}$. By Proposition 3.23, we conclude that $\mu=\mu\,\circ\,T^{-1}$, namely, $\mu\notin U$. Hence, $U=\bigcup_{j\in\mathbb{N}}\Phi_{j}$, establishing the claim. Next, note that $\{\varphi_{j}\}_{j\in\mathbb{N}}$ is a uniformly computable sequence of functions and $T$ is computable. Then by Corollary 3.20, the family $\{\mathcal{I}_{j}\}_{j\in\mathbb{N}}$ of integral operators $\mathcal{I}_{j}\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ defined by
$\mathcal{I}_{j}(\mu)\coloneqq\int\!(\varphi_{j}-\varphi_{j}\circ T)\,\mathrm{d}\mu$
is a uniformly computable sequence of functions. By Definition 3.5, $\mathbb{R}\setminus\{0\}$ is a lower-computable open set. Note that $\Phi_{j}=\mathcal{I}_{j}^{-1}(\mathbb{R}\setminus\{0\})$. Then by Corollary 3.11, $\{\Phi_{j}\}_{j\in\mathbb{N}}$ is a uniformly computable sequence of lower-computable open sets. Therefore, by Proposition 3.7 and the claim, $U$ is a lower-computable open set. ∎
Lemma 4.13.
Let the quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfy the Assumption A in Section 1. Define the set $W$ so that a Borel probability measure $\mu$ is in $W$ if and only if the Jacobian of $T\colon X\rightarrow X$ with respect to $\mu$ is not
$J_{\varphi}(x)\coloneqq\frac{u_{\varphi}(T(x))}{u_{\varphi}(x)}\exp(P(T,\varphi)-\varphi(x)).$
Then $W$ is a lower-computable open set.
Proof.
Let $\{\varphi_{j}\}_{j\in\mathbb{N}}$ be a computable enumeration of $\mathcal{E}(\mathcal{S})$ (see Definition 3.22) in the computable metric space $(X,\,\rho,\,\mathcal{S})$, and $\{x_{k}:1\leqslant k\leqslant m\}$ be an $\eta$-net for $X$. Set $g_{k,n}(x)\coloneqq g_{x_{k},\eta-\frac{1}{n},\frac{1}{n}}(x)$ (see Definition 3.21) for each $n\in\mathbb{N}$ and each integer $k\leqslant m$. Claim. $W=\bigcup_{k=1}^{m}\bigcup_{j,\,n\in\mathbb{N}}\Phi_{j}^{k,n}$, where
(4.21) $\Phi_{j}^{k,n}\coloneqq\left\{\mu\in\mathcal{P}(X):\int\!(\varphi_{j}\cdot g_{k,n})\circ T^{-1}\,\mathrm{d}\mu-\int\!\varphi_{j}\cdot g_{k,n}\cdot J_{\varphi}\,\mathrm{d}\mu\neq 0\right\}$
for all $j,\,k,\,n\in\mathbb{N}$ with $k\leqslant m$. By the definition of $W$ and Jacobians, $\Phi_{j}^{k,n}\subseteq W$ for all $j,\,k,\,n\in\mathbb{N}$ with $k\leqslant m$. Therefore, to prove the claim, it suffices to show $W\subseteq\bigcup_{k=1}^{m}\bigcup_{j,\,n\in\mathbb{N}}\Phi_{j}^{k,n}$. Fix an integer $1\leqslant k\leqslant m$. Since $T$ is distance-expanding with constants $\eta$, $\lambda,$ and $\xi$, it is injective on $B(x_{k},\eta)$. Thus, the function $(\varphi_{j}\cdot g_{k,n})\circ T^{-1}$ is well-defined in $X$ for all $j,\,n\in\mathbb{N}$. Consider
$\mu\in\bigcap\limits_{j,\,n\in\mathbb{N}}\bigl{(}\Phi^{k,n}_{j}\bigr{)}^{c}.$
Then for each $j\in\mathbb{N}$, by (4.21), we have
$\int\!(\varphi_{j}\cdot g_{k,n})\circ T^{-1}\,\mathrm{d}\mu=\int\!\varphi_{j}\cdot g_{k,n}\cdot J_{\varphi}\,\mathrm{d}\mu\quad\text{for each }n\in\mathbb{N}.$
Letting $n\to+\infty$, it follows from Dominated Covergence Theorem that
$\int\!\bigl{(}\varphi_{j}\cdot\mathbbm{1}_{B(x_{k},\eta)}\bigr{)}\circ T^{-1}\,\mathrm{d}\mu=\int\!\varphi_{j}\cdot\mathbbm{1}_{B(x_{k},\eta)}\cdot J_{\varphi}\,\mathrm{d}\mu$
for each $j\in\mathbb{N}$. Then by Proposition 3.23, it follows that $\mu_{1}=\mu_{2}$, where $\mu_{1}$ and $\mu_{2}$ are given by
$\mu_{1}(A)\coloneqq\frac{\mu(T(A\cap B(x_{k},\eta)))}{\mu(T(B(x_{k},\eta)))}\quad\text{and}\quad\mu_{2}(A)\coloneqq\frac{\int_{A\cap B(x_{k},\eta)}\!J_{\varphi}\,\mathrm{d}\mu}{\int_{B(x_{k},\eta)}\!J_{\varphi}\,\mathrm{d}\mu}$
for each Borel set $A\subseteq X$. Thus,
$\mu(T(A\cap B(x_{k},\eta)))=\int_{A\cap B(x_{k},\eta)}\!J_{\varphi}\,\mathrm{d}\mu$
for each Borel set $A\subseteq X$. Therefore, for each $\mu\in\bigcap\limits_{1\leqslant k\leqslant m}\bigcap\limits_{j,\,n\in\mathbb{N}}\bigl{(}\Phi^{k,n}_{j}\bigr{)}^{c}$ and Borel set $A\subseteq X$, we can define $A_{1}\coloneqq A\cap B(x_{1},\eta)$, and $A_{k}\coloneqq\bigl{(}A-\bigcup_{i=1}^{k-1}A_{i}\bigr{)}\cap B(x_{k},\eta)$ for each integer $k\in[2,m]$. Since $X=\bigcup_{k=1}^{m}B(x_{k},\eta)$, then $A=\bigsqcup_{k=1}^{m}A_{k}$. Hence,
$\mu(T(A))=\sum\limits_{k=1}^{m}\mu(T(A_{k}))=\sum\limits_{k=1}^{m}\int_{A_{k}}\!J_{\varphi}\,\mathrm{d}\mu=\int_{A}\!J_{\varphi}\,\mathrm{d}\mu,$
namely, $\mu\in W^{c}.$ Hence, the claim holds. Now we return to the proof of Lemma 4.13. By Definition 3.21, Theorem 4.7, and Proposition 4.8, $\bigl{\{}(\varphi_{j}\cdot g_{k,n})\circ T^{-1}-\varphi_{j}\cdot g_{k,n}\cdot J_{\varphi}:j,\,k,\,n\in\mathbb{N}\text{ with }k\leqslant m\bigr{\}}$ is a uniformly computable sequence of functions. Then by Corollary 3.20, the family $\big{\{}\mathcal{I}_{j}^{k,n}:j,\,k,\,n\in\mathbb{N}\text{ with }k\leqslant m\big{\}}$ of integral operators $\mathcal{I}_{j}^{k,n}\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ defined by
$\mathcal{I}_{j}^{k,n}(\mu)\coloneqq\int\!\bigl{(}(\varphi_{j}\cdot g_{k,n})\circ T^{-1}-\varphi_{j}\cdot g_{k,n}\cdot J_{\varphi}\bigr{)}\,\mathrm{d}\mu,\qquad\mu\in\mathcal{P}(X),$
is a uniformly computable sequence of functions. By Definition 3.5, $\mathbb{R}\setminus\{0\}$ is a lower-computable open set. Note that $\Phi_{j}^{k,n}=\bigl{(}\mathcal{I}_{j}^{k,n}\bigr{)}^{-1}(\mathbb{R}\setminus\{0\})$. Then by Corollary 3.11, $\Phi_{j}^{k,n}:j,\,k,\,n\in\mathbb{N}\text{ with }k\leqslant m$ is a uniformly computable sequence of lower-computable open sets. Therefore, by Proposition 3.7 and the claim, $W$ is a lower-computable open set. ∎
4.4. Proofs of Theorems 1.1 and 1.2
We are now ready to complete the proof of Theorem 1.1.
Proof of Theorem 1.1.
Assume that $X$ is recursively compact, then by Proposition 3.18, $\mathcal{P}(X)$ is recursively compact. By Corollary 2.9, there exists a unique $T$-invariant Gibbs state $\mu_{\varphi}$ for $T$ and $\varphi$, hence by Proposition 2.5, an unique equilibrium state $\mu_{\varphi}$ of $T$ for $\varphi$. Together with the definitions of $U$ and $W$, we have $\{\mu_{\varphi}\}=\mathcal{P}(X)\setminus(U\cup W)$. By Lemmas 4.12 and 4.13, $U$ and $W$ are both lower-computable open sets. It follows from Proposition 3.15 (ii) that $\{\mu_{\varphi}\}$ is recursively compact. Therefore, by Proposition 3.15 (i), the equilibrium state $\mu_{\varphi}$ of $T$ for $\varphi$ is computable. ∎
Applying Theorem 1.1, we establish Theorem 1.2.
Proof of Theorem 1.2.
In the computable metric space $(D,\,\rho,\,\mathcal{S}^{\prime})$, by Theorem 1.1, there exists an algorithm which, on input $n\in\mathbb{N}$, outputs $\{k_{i}\}_{i=1}^{m}\subseteq\mathbb{Q}^{+}$ and $\{x_{i}\}_{i=1}^{m}\subseteq\mathcal{S}^{\prime}$ satisfying that $\sum_{i=1}^{m}k_{i}=1$, and $W_{\rho}(\mu_{\varphi},\mu_{n})\leqslant 2^{-n},$ where $\mu_{n}=\sum_{i=1}^{m}k_{i}\delta_{x_{i}}\in\mathcal{R}_{\mathcal{S}^{\prime}}$. Note that $\mathcal{S}^{\prime}$ is a uniformly computable sequence of points in the computable metric space $(X,\,\rho,\,\mathcal{S})$. So for each integer $1\leqslant i\leqslant m$, we can find a point $y_{i}\in\mathcal{S}$ such that $\rho\bigl{(}x_{i},y_{i}\bigr{)}<2^{-n}$. Define $\widehat{\mu}_{n}\coloneqq\sum_{i=1}^{m}k_{i}\delta_{y_{i}}\in\mathcal{R}_{\mathcal{S}}.$ Then
$W_{\rho}(\mu_{\varphi},\widehat{\mu}_{n})\leqslant W_{\rho}(\mu_{\varphi},\mu_{n})+W_{\rho}(\mu_{n},\widehat{\mu}_{n})<2^{1-n}.$
Therefore, $\mu_{\varphi}$ is a computable point in the computable metric space $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{R}_{\mathcal{S}})$. ∎
5. Hyperbolic rational maps
In this section, we review some notations and results in complex dynamics and apply Theorem 1.2 to prove Theorem C. Recall that the spherical metric $d_{\widehat{\mathbb{C}}}$ on the Riemann sphere ${}^\widehat{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$, given by the length element $\mathrm{d}s(z)\coloneqq 2\lvert\mathrm{d}z\rvert/(1+\lvert z\rvert^{2})$ for each $z\in\widehat{\mathbb{C}}$, is a conformal metric. Moreover, (^$\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$ is a computable metric space, where $\mathbb{Q}^{2}$ is defined by $\{a+bi:a,\,b\in\mathbb{Q}\}$, and $\widehat{\mathbb{C}}$ is recursively compact. The spherical derivative $f^{\#}$ of a holomorphic function $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is given by
(5.1) $f^{\#}(z)\coloneqq\lim\limits_{w\to z}\frac{d_{\widehat{\mathbb{C}}}(f(w),f(z))}{d_{\widehat{\mathbb{C}}}(w,z)}=\frac{1+\lvert z\rvert^{2}}{1+\lvert f(z)\rvert^{2}}\cdot\lvert f^{\prime}(z)\rvert$
for each $z\in\widehat{\mathbb{C}}$. If $z$ or $f(z)$ is equal to $\infty$, then the last expression of (5.1) has to be understood as a suitable limit. We identify $\widehat{\mathbb{C}}$ with the unit sphere in $\mathbb{R}^{3}$ via stereographic projection. The chordal metric $\sigma$ on $\widehat{\mathbb{C}}$ is the metric that corresponds to the Euclidean metric in $\mathbb{R}^{3}$ under this identification. More precisely, we have
(5.2) $\sigma(z,w)=\frac{2\lvert z-w\rvert}{\sqrt{1+\lvert z\rvert^{2}}\sqrt{1+\lvert w\rvert^{2}}}\quad\text{for each }z,\,w\in\mathbb{C},$
and $\sigma(\infty,z)=\sigma(z,\infty)=2\big{/}\sqrt{1+\lvert z\rvert^{2}}$ for each $z\in\mathbb{C}$. Any holomorphic map $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ on the Riemann sphere can be expressed as a rational function, that is, as the quotient $f(z)=p(z)/{q(z)}$ of two polynomials. Here we may assume that $p(z)$ and $q(z)$ have no common roots. The degree $d$ of the rational map $f$ is defined as the maximum of the degrees of $p$ and $q$. The Fatou set $\mathcal{F}_{f}$ of $f$ is defined to be the set of points $z\in\widehat{\mathbb{C}}$ such that there exists an open neighborhood $U(z)$ on which the family $\{f^{n}|_{U(z)}\}_{n\in\mathbb{N}}$ is equicontinuous with respect to the spherical metric $d_{\widehat{\mathbb{C}}}$. The Julia set $\mathcal{J}_{f}$ of $f$ is the complement of the Fatou set $\mathcal{F}_{f}$.
Proposition 5.1 ([Mil06, Corollary 4.13]).
If $-^f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is a rational map whose degree $d\geqslant 2$, then for each $z_{0}\in\mathcal{J}_{f}$, the set $\bigcup_{n\in\mathbb{N}}f^{-n}(z_{0})$ of all iterated preimages of $z_{0}$ is dense in $\mathcal{J}_{f}$.
The map we consider in this section is a hyperbolic rational map, which is defined as follows (see for example, [Mil06, Theorem 19.1]).
Definition 5.2.
A rational function $f$ of degree $d\geqslant 2$ is hyperbolic if the closure of the set of postcritical points of $f$ is disjoint from $\mathcal{J}_{f}$.
Recall that a point $x\in\widehat{\mathbb{C}}$ is a postcritical point of $f$ if $x=f^{n}(c)$ for some critical point $c\in\widehat{\mathbb{C}}$ of $f$ and some $n\in\mathbb{N}$. For hyperbolic rational maps, we have the following two results.
Lemma 5.3 ([CG93, Lemma 2.1]).
Assume that $-^f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is a hyperbolic rational map whose degree is at least two with $\infty\notin\mathcal{J}_{f}$. Then there exist two constants $a>0$ and $A>1$ for all $n\in\mathbb{N}$ and $z\in\mathcal{J}_{f}$, $\lvert(f^{n})^{\prime}(z)\rvert\geqslant aA^{n}$.
Lemma 5.4 ([CG93, Theorem 2.3]).
If $f$ is a hyperbolic rational map whose degree is at least two, then $\mathcal{J}_{f}$ has zero area.
The following lemma is due independently to Braverman [Br04] (in the case of polynomials) and Rettinger [Re04] (in the case of rational maps).
Lemma 5.5.
If $f$ is a computable hyperbolic rational map whose degree is at least two, then the distance function $d_{\widehat{\mathbb{C}}}(\mathcal{J}_{f},x)$ is computable.
Proposition 5.6.
Let $-^f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a computable hyperbolic rational map of degree $d\geqslant 2$. Then for each repelling periodic point $w$ of $f$, the following statements hold:
  1. (i)
    There is an enumeration $\{w_{i}\}_{i}$ of the set $\bigcup_{n\in\mathbb{N}}f^{-n}(w)$ such that $\{w_{i}\}_{i}$ is a uniformly computable sequence of points in the computable metric space (^$\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$;
  2. (ii)
    $\bigl{(}\mathcal{J}_{f},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}f^{-n}(w)\bigr{)}$ is a computable metric space, and the Julia set $\mathcal{J}_{f}$ is recursively compact in the computable metric space $\bigl{(}\mathcal{J}_{f},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}f^{-n}(w)\bigr{)}$.
Proof.
(i) First, we demonstrate that every periodic point of $f$ is computable in the computable metric space (^$\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$. Let $u$ be a such point, then $f^{m}(u)=u$ for some $m\in\mathbb{N}$. To establish this, we apply a standard root-finding algorithm for the function $f^{m}(x)-x$ (see for example, [BBY10, Appendix A]) to compute $u$. Hence, $w$ is computable. By [Mil06, Lemma 4.6], we obtain that $w\in\mathcal{J}_{f}$. Hence, by the same root-finding algorithm, we can compute all the roots of the function $g_{n}(x)=f^{n}(x)-w$ for each $n\in\mathbb{N}$ and enumerate them as $\{w_{i}\}_{i}$. Hence, $\{w_{i}\}_{i}$ is a uniformly computable sequence of points in the computable metric space (^$\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$. (ii) By Proposition 5.1, $\bigcup_{n\in\mathbb{N}}f^{-n}(w)$ is dense in $\mathcal{J}_{f}$. Together with statement (i), one can see that $\bigl{(}\mathcal{J}_{f},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}f^{-n}(w)\bigr{)}$ is a computable metric space. Hence, to demonstrate statement (ii), by Proposition 4 in [GHR10], it suffices to prove that $\bigl{(}\mathcal{J}_{f},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}f^{-n}(w)\bigr{)}$ is recursively precompact. Fix an integer $m$. Since (^$\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$ is recursively precompact, we can compute a $2^{-m-1}$-net $N^{\prime\prime}_{m}$ of $\widehat{\mathbb{C}}$. By Lemma 5.5, we can compute a $2^{-n-1}$-net $N^{\prime}_{m}\subseteq N^{\prime\prime}_{m}$ of $\mathcal{J}_{f}$ satisfying that $\mathcal{J}_{f}\cap B(y,2^{-m-1})\neq\emptyset$ for each $y\in N^{\prime}_{m}$. By Proposition 5.1 and statement (i), we can compute a point $t(y)\in B(y,2^{-m-1})\cap\bigl{(}\bigcup_{n\in\mathbb{N}}f^{-n}(w)\bigr{)}$ for each $y\in N^{\prime}_{m}$. Hence, the set $\{t(y):y\in N^{\prime}_{m}\}\subseteq\mathcal{J}_{f}$ is a $2^{-m}$-net of the Julia set $\mathcal{J}_{f}$. This completes the proof. ∎
Proposition 5.7.
Let $-^f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a hyperbolic rational map of degree $d\geqslant 2$ with $\infty\notin\mathcal{J}_{f}$. Let $C_{f}\coloneqq f(\{z\in\widehat{\mathbb{C}}:f^{\prime}(z)=0\})$ be the $f$-image of the critical set of $f$. Assume that there exist constants $V_{1}$, $V_{2}$, $V_{3}$, and $V_{4}$ satisfying the following properties for each pair $x,\,y\in\mathcal{J}_{f}$:
  1. (1)
    $B(x,V_{1})$ is disjoint from $C_{f}$;
  2. (2)
    $8/V_{4}<V_{2}\leqslant\lvert f^{\prime}(x)\rvert\leqslant V_{3}$;
  3. (3)
    $d_{\widehat{\mathbb{C}}}(x,y)\geqslant V_{4}\lvert x-y\rvert$.
Then the following statements hold:
  1. (i)
    $f$ is injective on $B(x,l)$ for each $x\in\mathcal{J}_{f}$, where $l\coloneqq V_{1}/(4V_{2})$.
  2. (ii)
    $f$ is distance-expanding on $\mathcal{J}_{f}$ with respect to the spherical metric with constants $\eta\coloneqq V_{4}l$, $\lambda\coloneqq V_{2}V_{4}/8$, and $\xi\coloneqq V_{2}l/8$, where $l$ is from statement (i).
Proof.
(i) Fix an arbitrary point $x\in\mathcal{J}_{f}$. By property (1), one can see that there exists no critical value in $B(f(x),V_{1})$. Hence, we can define a conformal inverse $f^{-1}_{x}\colon B(f(x),V_{1})\rightarrow\widehat{\mathbb{C}}$ in the sense of $f(f^{-1}_{x}(y))=y$ for each $y\in B(f(x),V_{1})$ and $f^{-1}_{x}(f(x))=x$. Then by Köbe’s distorsion theorem, we have $B(x,\lvert(f_{x}^{-1})^{\prime}(f(x))\rvert V_{1}/4)\subseteq f^{-1}_{x}(B(f(x),V_{1}))$. By the Chain Rule and property (2), $\lvert(f_{x}^{-1})^{\prime}(f(x))\rvert=1/\lvert f^{\prime}(x)\rvert\geqslant V_{3}^{-1}$. Thus, one can see that $l=V_{1}/(4V_{3})\leqslant\lvert(f_{x}^{-1})^{\prime}(f(x))\rvert V_{1}/4$. Therefore, $f$ is injective on $B(x,l)$. (ii) Fix an arbitrary pair of $x,\,y\in\mathcal{J}_{f}$ with $d_{\widehat{\mathbb{C}}}(x,y)\leqslant\eta$. First, we demonstrate that $d_{\widehat{\mathbb{C}}}(f(x),f(y))\geqslant\lambda d_{\widehat{\mathbb{C}}}(x,y)$. By property (3), we obtain that $\lvert x-y\rvert\leqslant\eta/V_{4}=l$. Hence, by statement (i), $f$ is injective on $B(x,\lvert x-y\rvert)$. By Köbe’s distorsion theorem, we have
$B(f(x),\lvert x-y\rvert\cdot\lvert f^{\prime}(x)\rvert/4)\subseteq f(B(x,\lvert x-y\rvert)).$
Noting that $f(y)\in\partial f(B(x,\lvert x-y\rvert))$, one can see that $\lvert f(x)-f(y)\rvert\geqslant\lvert x-y\rvert\cdot\lvert f^{\prime}(x)\rvert/4\geqslant V_{2}\lvert x-y\rvert/4$. Since $\mathcal{J}_{f}$ is $f$-invariant, then $f(x),\,f(y)\in\mathcal{J}_{f}$. By property (3), we have
$d_{\widehat{\mathbb{C}}}(f(x),f(y))\geqslant V_{4}\lvert f(x)-f(y)\rvert\geqslant V_{2}V_{4}\lvert x-y\rvert/4\geqslant V_{2}V_{4}d_{\widehat{\mathbb{C}}}(x,y)/8=\lambda d_{\widehat{\mathbb{C}}}(x,y).$
Next, we verify that $B_{d_{\widehat{\mathbb{C}}}}(f(x),\xi)\subseteq f(B_{d_{\widehat{\mathbb{C}}}}(x,\eta))$. To see this, since $V_{4}\leqslant\frac{d_{\widehat{\mathbb{C}}}(x,y)}{\lvert x-y\rvert}\leqslant 2$, it suffices to check that $B(f(x),2\xi)\subseteq f(B(x,l))$. By statement (i), it follows from by Köbe’s distorsion theorem that
$B(f(x),2\xi)=B(f(x),V_{2}l/4)\subseteq B(f(x),\lvert f^{\prime}(x)\rvert l/4)\subseteq f(B(x,l)).$
This completes the proof of statement (ii). ∎
Proposition 5.8.
Let $-^f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a hyperbolic rational map of degree $d\geqslant 2$ with $\infty\notin\mathcal{J}_{f}$. Let $x_{k}$, $1\leqslant k\leqslant m$, be an enumeration of $f^{-1}(\infty)\cup\{\infty\}$, and $r>0$ be a constant satisfying that $\mathcal{J}_{f}\subseteq K\coloneqq\bigcap_{k=1}^{m}B_{d_{\widehat{\mathbb{C}}}}(x_{k},r)^{c}$ and $B_{d_{\widehat{\mathbb{C}}}}(x_{i},r)\cap B_{d_{\widehat{\mathbb{C}}}}(x_{j},r)=\emptyset$ for all $1\leqslant i<j\leqslant m$. Fix constants $D_{1},\,D_{2},\,C_{1},\,C_{2},\,C>0$ satisfy the following conditions:
  1. (1)
    $\lvert z\rvert\leqslant D_{1}$ and $\lvert f(z)\rvert\leqslant D_{2}$ for each $z\in K$;
  2. (2)
    $\lvert f^{\prime}(z)\rvert\leqslant C_{1}$ and $\lvert f^{\prime\prime}(z)\rvert\leqslant C_{2}$ for each $z\in K$;
  3. (3)
    $\lvert f^{\#}(z)\rvert\geqslant C$ for each $z\in\mathcal{J}_{f}$.
Then the following statements hold:
  1. (i)
[…]
    $d_{K}(x,y)\coloneqq\inf\biggl{\{}\int_{\gamma}\!\mathrm{d}s:\gamma(0)=x,\,\gamma(1)=y,\,\text{and }\gamma\colon[0,1]\rightarrow K\text{ is continuous}\biggr{\}};$
  2. (ii)
    If $g\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is a rational map, then $\lvert g\rvert$ is $\bigl{(}A(1+D_{1}^{2})\pi/2\bigr{)}$-Lipschitz continuous with respect to $d_{\widehat{\mathbb{C}}}$ on $K$, where $A\coloneqq\sup_{z\in K}\lvert g^{\prime}(z)\rvert$;
  3. (iii)
    $\varphi_{f}(z)$ is $(BC^{-1})$-Lipschitz continuous on $\mathcal{J}_{f}$ with respect to $d_{\widehat{\mathbb{C}}}$, where $\varphi_{f}(z)\coloneqq\log\bigl{(}f^{\#}(z)\bigr{)}$ for each $z\in\mathcal{J}_{f}$, where $B\coloneqq(1+D_{1}^{2})\bigl{(}C_{2}(1+D_{1}^{2})+D_{2}C_{1}^{2}(1+D_{1}^{2})+C_{1}D_{1}\bigr{)}\pi$.
Proof.
(i) The path-connectedness of $K$ is immediate from its definition. Hence, $d_{K}$ is well-defined. By subdividing the spherical geodesic from $x$ to $y$, we only need to consider the case where $x,\,y\in\partial B_{d_{\widehat{\mathbb{C}}}}(x_{i},r)$ for some $1\leqslant i\leqslant m$. Then we have
$d_{K}(x,y)\leqslant\pi\sigma(x,y)\leqslant\pi d_{\widehat{\mathbb{C}}}(x,y).$
(ii) Fix an arbitrary pair of $x,\,y\in K$. Since $K$ is closed, there exists a continuous function $\gamma\colon[0,1]\mapsto K$ with $\gamma(0)=x,\,\gamma(1)=y,$ and $d_{K}(x,y)=\int_{\gamma}\!\mathrm{d}s$. Hence, for each pair of $x,\,y\in K$, we have
$\lvert g(x)-g(y)\rvert=\biggl{\lvert}\int_{\gamma}\!g^{\prime}(z)\,\mathrm{d}z\biggr{\rvert}\leqslant\int_{\gamma}\!\frac{A(1+\lvert z\rvert^{2})}{2}\,\mathrm{d}s(z)\leqslant\frac{A(1+D_{1}^{2})d_{K}(x,y)}{2}\leqslant\frac{A(1+D_{1}^{2})\pi d_{\widehat{\mathbb{C}}}(x,y)}{2}.$
(iii) Define $u(x)=1+\lvert x\rvert^{2}$, $v(x)\coloneqq\frac{1}{1+\lvert f(x)\rvert^{2}}$, and $w(x)\coloneqq\lvert f^{\prime}(x)\rvert$ for each $x\in K$. Then we establish that $u,v,$ and $w$ are all Lipschitz continuous functions, respectively. By statement (i), conditions (1) and (2), we have
(5.3) $\lvert u(x)-u(y)\rvert=(\lvert x\rvert+\lvert y\rvert)\left\lvert\lvert x\rvert-\lvert y\rvert\right\rvert\leqslant 2D_{1}\lvert x-y\rvert\leqslant D_{1}(1+D_{1}^{2})\pi d_{\widehat{\mathbb{C}}}(x,y),$
(5.4) $\lvert v(x)-v(y)\rvert=\frac{\lvert\lvert f(x)\rvert^{2}-\lvert f(y)\rvert^{2}\rvert}{(1+\lvert f(x)\rvert^{2})(1+\lvert f(y)\rvert^{2})}\leqslant 2D_{2}\lvert\lvert f(x)\rvert-\lvert f(y)\rvert\rvert\leqslant D_{2}C_{1}(1+D_{1}^{2})\pi d_{\widehat{\mathbb{C}}}(x,y),$
and
(5.5) $\lvert w(x)-w(y)\rvert\leqslant C_{2}(1+D_{1}^{2})\pi d_{\widehat{\mathbb{C}}}(x,y)$
for each pair of $x,\,y\in K$. By the definition of $K$, (5.1) is true for each $z\in K$. Moreover, by (5.3), (5.4), and (5.5), we have
(5.6) $\bigl{\lvert}f^{\#}(x)-f^{\#}(y)\bigr{\rvert}$
$\qquad\leqslant\lvert u(x)v(x)\rvert\lvert w(x)-w(y)\rvert+\lvert u(x)w(y)\rvert\lvert v(x)-v(y)\rvert+\lvert v(y)w(y)\rvert\lvert u(x)-u(y)\rvert$
$\qquad\leqslant\bigl{(}C_{2}(1+D_{1}^{2})^{2}+D_{2}C_{1}^{2}(1+D_{1}^{2})^{2}+C_{1}D_{1}(1+D_{1}^{2})\bigr{)}\pi d_{\widehat{\mathbb{C}}}(x,y)$
$\qquad=Bd_{\widehat{\mathbb{C}}}(x,y)$
for each pair of $x,\,y\in K$. Note that $\log x<x-1$ for each $x>1$. Therefore, it follow from condition (3) that
$\lvert\varphi_{f}(x)-\varphi_{f}(y)\rvert=\biggl{\lvert}\log\biggl{(}\frac{\lvert f^{\#}(x)\rvert}{\lvert f^{\#}(y)\rvert}\biggr{)}\biggr{\rvert}\leqslant\frac{\lvert f^{\#}(x)-f^{\#}(y)\rvert}{\min\{\lvert f^{\#}(x)\rvert,\,\lvert f^{\#}(y)\rvert\}}\leqslant BC^{-1}d_{\widehat{\mathbb{C}}}(x,y)$
for each pair of $x,\,y\in\mathcal{J}_{f}$. ∎
Now we are ready to establish Theorem C.
Proof of Theorem C.
Since $f$ is a hyperbolic rational map, it follows from Lemma 5.4 that there exists a computable point $u$ in $\mathcal{J}_{f}^{c}$. By Lemma 5.5, we can compute a point $u\in\mathcal{J}_{f}^{c}$. Define a new function $g$ by
$g(z)\coloneqq\bigl{(}U\circ f\circ U^{-1}\bigr{)}(z)=1\big{/}\bigl{(}f\bigl{(}u+z^{-1}\bigr{)}-u\bigr{)}$
for each $z\in\widehat{\mathbb{C}}$. It follows from the computability of $f$ and $u$ that $g$ is a computable hyperbolic rational map with $\infty\notin\mathcal{J}_{g}$. First, we compute constants $n,\,\eta,\,\lambda$, and $\xi$ which satisfy that $g^{n}$ is distance-expanding on $\mathcal{J}_{g}$ with respect to the spherical metric with $\eta,\,\lambda$, and $\xi$. We apply a standard root-finding algorithm to compute the periodic points of $g$ as well as their multipliers. We can semi-decide whether a periodic point is repelling (if the multiplier is greater than $1$). Hence, we can compute the repelling periodic point $w$ of $g$. By Proposition 5.6, one can see that $\mathcal{J}_{g}$ is recursively compact in the computable metric space $\bigl{(}\mathcal{J}_{g},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}g^{-n}(w)\bigr{)}$. By [BRY12, Proposition 2.13], we can compute a constant $r_{0}>0$ with $\lvert z\rvert\leqslant r_{0}$ for each $z\in\mathcal{J}_{g}$. Then by (5.2), one can conclude that
$d_{\widehat{\mathbb{C}}}(x,y)\geqslant\sigma(x,y)\geqslant 2\lvert x-y\rvert\big{/}\bigl{(}1+r_{0}^{2}\bigr{)}$
[…]
6. Non-uniqueness of equilibrium states
In this section, we establish Theorem E and then use it to prove Theorem D. We first recall a classical result in Computable Analysis.
Lemma 6.1 ([GHR10, Proposition 7]).
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Assume that $X$ is recursively compact, and $\mu\in\mathcal{P}(X)$ is computable. Then there exists at least one computable point in the support of $\mu$.
Lemma 6.2.
Let $(X,\rho)$ be a compact metric space, and $T\colon X\rightarrow X$ be a continuous map. Then $\operatorname{supp}(\mu)\subseteq\Omega(T)$ for each $\mu\in\mathcal{M}_{T}(X)$.
$\Omega(T)$ is the non-wandering set of $T$ defined in (2.1).
Proof.
Consider $\mu\in\mathcal{M}_{T}(X)$. By [Wa82, Theorem 5.6 (i)], $\Omega(T)$ is closed. Moreover, it follows from [Wa82, Theorem 6.15 (i)] that $\mu(\Omega(T))=1$. Therefore, $\operatorname{supp}(\mu)\subseteq\Omega(T)$. ∎
Lemma 6.3.
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space and $X$ be a recursively compact set. Assume that $T\colon X\rightarrow X$ is a computable map satisfying that the measure-theoretic entropy function $\mathcal{H}\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ given by
$\mathcal{H}(\mu)\coloneqq\begin{cases}h_{\mu}(T)&\text{if }\mu\in\mathcal{M}_{T}(X);\\ 0&\text{if }\mu\notin\mathcal{M}_{T}(X)\end{cases}$
is upper-computable on $\mathcal{M}_{T}(X)$, and that $\varphi\colon X\rightarrow\mathbb{R}$ is a computable potential satisfying that the topological pressure $P(T,\varphi)$ is lower-computable. Then the set of equilibrium states of $T$ for $\varphi$ is recursively compact.
Proof.
By Corollary 3.20, the integral function $\Phi\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ given by $\Phi(\mu)\coloneqq\int\!\varphi\,\mathrm{d}\mu$, $\mu\in\mathcal{P}(X)$, is computable. Since $\mathcal{H}(\mu)$ is upper-computable on $\mathcal{M}_{T}(X)$, then the measure-theoretic pressure function $\mathcal{P}(\mu)\coloneqq\mathcal{H}(\mu)+\Phi(\mu)$ is upper-computable on $\mathcal{M}_{T}(X)$. Since $P(T,\varphi)$ is lower-computable, then there exists a uniformly computable sequence of rational points $\{r_{n}\}_{n\in\mathbb{N}}$ which is increasing and converges to $P(T,\varphi)$. By Definition 3.12, there exists a uniformly sequence $\{U_{n}\}_{n\in\mathbb{N}}$ of lower-computable open sets satisfying that $\mathcal{P}^{-1}(-\infty,r_{n})\cap\mathcal{M}_{T}(X)=U_{n}\cap\mathcal{M}_{T}(X)$ for each $n\in\mathbb{N}$. Hence,
(6.1) $\mathcal{M}_{T}(X)\setminus\mathcal{P}^{-1}(-\infty,P(T,\varphi))$ $=\mathcal{M}_{T}(X)\setminus\Bigl{(}\bigcup_{n\in\mathbb{N}}\mathcal{P}^{-1}(-\infty,r_{n})\cap\mathcal{M}_{T}(X)\Bigr{)}$
$=\Bigl{(}\bigcup_{n\in\mathbb{N}}U_{n}\Bigr{)}^{c}\cap\mathcal{M}_{T}(X).$
By Definition 2.4, this set consists of all equilibrium states of $T$ for $\varphi$. Moreover, by Propositions 3.7 and 3.15, it follows from (6.1) that this set is recursively compact. ∎
Now we are ready to prove Theorem E.
Proof of Theorem E.
We prove Theorem E by contradiction and suppose that $\mu$ is the unique equilibrium state of $T$ for $\varphi$. Then by Proposition 3.15 (i), it follows from Lemma 6.3 that $\mu$ is computable, which would contradict with the existence of non-computable equilibrium state. Noting that every upper-computable function is upper semi-continuous, one can see that $h_{\mu}(T)$ is upper semi-continuous on $\mu\in\mathcal{M}_{T}(X)$. Hence, by [PU10, Proposition 3.5.3], there exists at least one equilibrium state of $T$ for $\varphi$. Thus, statement (ii) holds. Therefore, there are at least two equilibrium states of $T$ for $\varphi$. ∎
As an immediate consequence, we establish Theorem D.
Proof of Theorem D.
Assume that $T\colon X\rightarrow X$ is a computable map with zero topological entropy satisfying that there exists no computable point in $\Omega(T)$. Then by [PU10, Proposition 3.5.3], there exists at least one equilibrium state of $T$ for $\varphi$. Suppose that $\mu$ is an equilibrium state of $T$ for $\varphi$. Then by Lemma 6.2, there exists no computable points in $\operatorname{supp}(\mu)$. It follows from Lemma 6.1 that $\mu$ is non-computable. Therefore, we can derive Theorem E from Theorem D. ∎
7. Counterexamples
7.1. Proof of Theorem F
In this subsection, we demonstrate that the computable homeomorphism $T\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ constructed in [GHR10, Subsection 4.1] satisfies the statements in Theorem F for each potential $\varphi\in C(\mathbb{S}^{1})$.
Proof of Theorem F.
Recall that $T$ constructed in [GHR10, Subsection 4.1] is a computable homeomorphism of the unit circle $\mathbb{S}^{1}$ whose non-wandering set $\Omega(T)$ contains no computable points. Since $T$ is a homeomorphism of the circle, it follows from [Wa82, Theorem 7.14] that the topological entropy of $T$ is zero, which implies that $T$ satisfies statement (i) of Theorem F. Suppose that $\mu$ is an equilibrium state of $T$ for $\varphi$. Then by Lemma 6.2, there exists no computable points in $\operatorname{supp}(\mu)$. It follows from Lemma 6.1 that $\mu$ is non-computable. Therefore $T$ satisfies statement (ii) of Theorem F. ∎
7.2. Asymptotic h-expansive
In this subsection, we introduce additional relevant notions and results from ergodic theory. We use them to construct computable systems with arbitrarily high topological entropy whose equilibrium states are all non-computable. We start by recalling the following definition from [Dow11, Remark 6.1.7].
Definition 7.1.
A sequence of open covers $\{\xi_{i}\}_{i\in\mathbb{N}_{0}}$ of a compact metric space $X$ is a refining sequence of open covers of $X$ if the following conditions are satisfied:
  1. (i)
    $\xi_{i+1}$ is a refinement of $\xi_{i}$ for each $i\in\mathbb{N}_{0}$.
  2. (ii)
    For each open cover $\eta$ of $X$, there exists $j\in\mathbb{N}$ such that $\xi_{i}$ is a refinement of $\eta$ for each integer $i\geqslant j$.
By the Lebesgue Covering Lemma, it is clear that for a compact metric space, refining sequences of open covers always exist. The topological tail entropy was first introduced by M. Misiurewicz under the name “topological conditional entropy” [Mis73, Mis76]. We use the terminology in [Dow11] (see [Dow11, Remark 6.3.18] for reference).
Definition 7.2.
Let $(X,d)$ be a compact metric space and $g\colon X\rightarrow X$ a continuous map. The topological conditional entropy $h(g|\lambda)$ of $g$ given $\lambda$, for some open cover $\lambda$, is defined as
(7.1) $h(g|\lambda)\coloneqq\lim\limits_{l\to+\infty}\lim\limits_{n\to+\infty}\frac{1}{n}H\bigl{(}(\xi_{l})_{g}^{n}\big{|}\lambda_{g}^{n}\bigr{)},$
where $\{\xi_{l}\}_{l\in\mathbb{N}_{0}}$ is an arbitrary refining sequence of open covers, and for each pair of open covers $\xi$ and $\eta$,
(7.2) $H(\xi|\eta)\coloneqq\log\Bigl{(}\max_{A\in\eta}\bigl{\{}\min\bigl{\{}\operatorname{card}\xi_{A}:\xi_{A}\subseteq\xi,\,A\subseteq\bigcup\xi_{A}\bigr{\}}\bigr{\}}\Bigr{)}$
is the logarithm of the minimal number of sets from $\xi$ sufficient to cover any set in $\eta$. The topological tail entropy $h^{*}(g)$ of $g$ is defined by
$h^{*}(g)\coloneqq\lim\limits_{m\to+\infty}\lim\limits_{l\to+\infty}\lim\limits_{n\to+\infty}\frac{1}{n}H\bigl{(}(\xi_{l})_{g}^{n}\big{|}(\eta_{m})_{g}^{n}\bigr{)},$
where $\{\xi_{l}\}_{l\in\mathbb{N}_{0}}$ and $\{\eta_{m}\}_{m\in\mathbb{N}_{0}}$ are two arbitrary refining sequences of open covers, and $H$ is as defined in (7.2).
The concept of $h$-expansiveness was introduced by R. Bowen in [Bo72]. We use the formulation in [Mis76] (see also [Dow11]).
Definition 7.3 ($h$-expansive).
A continuous map $g\colon X\rightarrow X$ on a compact metric space $(X,d)$ is called $h$-expansive if there exists a finite open cover $\lambda$ of $X$ such that $h(g|\lambda)=0$.
Remarks 7.4.
As an equivalent definition of $h$-expansiveness, the map $g\colon X\rightarrow X$ is called $h$-expansive if there exists $\epsilon>0$ such that
$h_{\operatorname{top}}\bigl{(}g|_{{\Phi}_{\epsilon,g}(x)}\bigr{)}=0$
for each $x\in X$, where
(7.3) $\Phi_{\epsilon,g}(x)\coloneqq\{y\in X:d(g^{n}(x),g^{n}(y))\leqslant\epsilon\text{ for each }n\in\mathbb{N}_{0}\}.$
A weaker property was introduced by M. Misiurewicz in [Mis73] (see also [Mis76, Dow11]).
Definition 7.5 (Asymptotic $h$-expansive).
We say that a continuous map $g\colon X\rightarrow X$ on a compact metric space $X$ is asymptotically $h$-expansive if $h^{*}(g)=0$.
Remarks 7.6.
The topological entropy of $g$ is $h_{\operatorname{top}}(g)=h(g|\{X\})$, where $\{X\}$ is the open cover of $X$ consisting of only one open set $X$. It also follows from Definition 7.1, for open covers $\iota$ and $\eta$ of $X$, one can see $h(g|\iota)\leqslant h(g|\eta)$ if $\iota$ is a refinement of $\eta$. Hence, we have $h^{*}(g)\leqslant h_{\operatorname{top}}(g)$.
We shall say that a pair $(X,f)$ is a cascade if $f\colon X\rightarrow X$ is a continuous map on the non-empty compact Hausdorff space $X$. Under the above notations, we have the following two results.
Proposition 7.7 ([Mis76, Theorem 3.2]).
Let $(X_{1},f_{1})$ and $(X_{2},f_{2})$ be two cascades then
$h^{*}(f_{1}\times f_{2})=h^{*}(f_{1})+h^{*}(f_{2}),$
where $f_{1}\times f_{2}\colon X_{1}\times X_{2}\rightarrow X_{1}\times X_{2}$ is given by
(7.4) $(f_{1}\times f_{2})(x_{1},x_{2})\coloneqq(f_{1}(x_{2}),f_{2}(x_{2}))\qquad\text{ for all }x_{1}\in X_{1}\text{ and }x_{2}\in X_{2}.$
Proposition 7.8 ([Mis76, Corollary 8.1]).
Let $(X,f)$ be a cascade. If $f$ is asymptotically $h$-expansive, then for each $\varphi\in C(X)$, there exists at least one equilibrium state of $f$.
Now we are ready to prove Proposition 7.9, a precise version of Theorem G. Here we identify $\mathbb{S}$ with $\mathbb{R}/[0,1]$.
Proposition 7.9.
Assume $d\in\mathbb{N}$ and let $T_{d}\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ be given by $T_{d}(x)\coloneqq dx\pmod{1}$ for $x\in\mathbb{S}$. Let $\widehat{T}\coloneqq T\times T_{d}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}\times\mathbb{S}^{1}$ be defined as in (7.4) and $\widehat{\varphi}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{R}$ be a continuous potential. Then statements (i) and (ii) in Theorem G hold for $\widehat{T}$ and $\widehat{\varphi}$, and $h_{\operatorname{top}}\bigl{(}\widehat{T}\bigr{)}=\log d$.
Proof.
First, we demonstrate that $\widehat{T}$ satisfies statement (i) in Theorem G. Note that for the map $T_{d}$ and sufficiently small $\epsilon>0$, the set $\Phi_{\epsilon,T_{d}}(x)=\{x\}$ for each $x\in\mathbb{S}^{1}$ (see (7.3)). Then by Remark 7.4, $T_{d}$ is $h$-expansive, and thus $h^{*}(T_{d})=0$. Noting that $h_{\operatorname{top}}(T)=0$, it follows from Remark 7.6 that $h^{*}(T)\leqslant h_{\operatorname{top}}(T)=0$, hence, is equal to 00. Hence by Proposition 7.7, $h^{*}\bigl{(}\widehat{T}\bigr{)}=h^{*}(T)+h^{*}(T_{d})=0,$ which implies that $\widehat{T}$ is asymptotically $h$-expansive. By Proposition 7.8, for each continuous potential $\widehat{\varphi}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{R}$, there exists at least one equilibrium state of $\widehat{T}$. Finally, we argue that there exists no computable $\widehat{T}$-invariant Borel probability measure by contradiction and assume that $\widehat{\mu}\in\mathcal{M}_{\widehat{T}}(\mathbb{S}^{1}\times\mathbb{S}^{1})$ is computable. Then $\mu$ given by $\mu(A)\coloneqq\widehat{\mu}(A\times\mathbb{S}^{1})$ for each Borel set $A\subseteq\mathbb{S}^{1}$ is a computable $T$-invariant probability measure. This is impossible by Theorem F (ii). Thus $\widehat{T}$ satisfies statement (ii) in Theorem G. ∎